### v1

# RGNMR: A Gauss-Newton method for robust matrix completion with theoretical guarantees

Abstract
Recovering a low rank matrix from a subset of its entries, some of which may be corrupted, is known as the robust matrix completion (RMC) problem. Existing RMC methods have several limitations: they require a relatively large number of observed entries; they may fail under overparametrization, when their assumed rank is higher than the correct one; and many of them fail to recover even mildly ill-conditioned matrices. In this paper we propose a novel RMC method, denoted RGNMR, which overcomes these limitations. RGNMR is a simple factorization-based iterative algorithm, which combines a Gauss–Newton linearization with removal of entries suspected to be outliers. On the theoretical front, we prove that under suitable assumptions, RGNMR is guaranteed exact recovery of the underlying low rank matrix. Our theoretical results improve upon the best currently known for factorization-based methods. On the empirical front, we show via several simulations the advantages of RGNMR over existing RMC methods, and in particular its ability to handle a small number of observed entries, overparameterization of the rank and ill-conditioned matrices.
1 Introduction
Figure 1: Performance of RMC methods as a function of the number of observed entries measured by the oversampling ratio $\frac{|\Omega|}{r\cdot(n_{1}+n_{2}-r)}$ where $r\cdot(n_{1}+n_{2}-r)$ is the number of degrees of freedom of a rank $r$ matrix. (left) Median rel-RMSE; (right) Failure probability (±$\pm 1.96$ SE). The underlying matrix $L^{*}$ has a condition number $\kappa=2$.
Figure 2: Performance of RMC methods under overparameterization with input rank of $5+i$ for $i\in[0,5]$. (left) Median rel-RMSE; (right) Failure probability (±$\pm 1.96$ SE). The matrix $L^{*}$ has a condition number $\kappa=2$ and the oversampling ratio is $\frac{|\Omega|}{r\cdot(n_{1}+n_{2}-r)}=12$.
Figure 3: Performance of RMC methods as a function of the condition number. (left) Median rel-RMSE; (right) Failure probability (±$\pm 1.96$ SE). The oversampling ratio is $\frac{|\Omega|}{r\cdot(n_{1}+n_{2}-r)}=12$.
Low-rank matrices play a fundamental role in multiple scientific disciplines. As reviewed in Davenport and Romberg, (2016), in various applications, there is a need to recover a low rank matrix from only a subset of its entries. Examples include recommendation systems (Bennett et al.,, 2007), various problems in computer vision (Kennedy et al.,, 2016; Tulyakov et al.,, 2016; Miao and Kou,, 2021), and sensor network localization (Drineas et al.,, 2006). A key challenge in these and other applications is that some of the observed entries may be arbitrarily corrupted outliers.
In this work, we consider the robust matrix completion (RMC) problem, of recovering a low rank matrix from a subset of its entries, out of which a few are outliers. Formally, let $X\in\mathbb{R}^{n_{1}\times n_{2}}$ be a matrix with a decomposition $X=L^{*}+S^{*}$, where $L^{*}$ is a rank $r$ matrix and $S^{*}$ is a sparse corruption matrix whose few non-zero entries are arbitrary. The RMC problem is to recover $L^{*}$ from the subset $\{X_{i,j}\,|\,(i,j)\in\Omega\}$, where $\Omega\subset[n_{1}]\times[n_{2}]$ denotes the subset of observed entries in $X$.
Related work.
Over the past decade, the RMC problem was studied from both mathematical, computational and statistical perspectives. Several RMC algorithms have been proposed. In general, to solve the RMC problem the following optimization issues need to be addressed: (i) either promote a low rank solution or strictly enforce it; and (ii) suppress the potentially detrimental effect of the outliers, whose locations are a-priori unknown.
Regarding the first issue, one approach is to incorporate into the optimization objective a low-rank promoting penalty such as the nuclear norm (Candès et al.,, 2011; Nie et al.,, 2012; Klopp et al.,, 2017; Wong and Lee,, 2017; Chen et al.,, 2021). Under suitable assumptions on the matrices $L^{*}$, $S^{*}$ and on the set $\Omega$, these methods enjoy strong theoretical guarantees. However, these schemes are in general computationally slow, and do not scale well to large matrices. A different approach is to strictly enforce a rank $r$ solution. This can be done by optimization over the manifold of rank $r$ matrices (Yan et al.,, 2013; Cambier and Absil,, 2016), or by projecting matrices onto it. Other works factorize the target matrix as $L=UV^{\top}$, and optimize over the two factor matrices $U\in\mathbb{R}^{n_{1}\times r}$ and $V\in\mathbb{R}^{n_{2}\times r}$ (Yi et al.,, 2016; Lin et al.,, 2017; Zeng and So,, 2017; Huang et al.,, 2021; Wang et al.,, 2023). The resulting problem involves only $(n_{1}+n_{2})r$ variables instead of the $n_{1}\times n_{2}$ entries in $L$. Factorization based methods are thus in general more scalable and able to handle larger matrices.
Regarding the second issue, some works attenuate the effect of the outliers either by estimating the locations of the corrupted entries and removing them, or by estimating their values $S_{ij}$ for $(i,j)\in\Omega$ (Yan et al.,, 2013; Cherapanamjeri et al.,, 2017; Huang et al.,, 2021; Chen et al.,, 2021; Wang and Wei,, 2024). Other works mitigate the influence of outliers by using a robust norm on the difference $X-L$ in the objective function (Nie et al.,, 2012; Cambier and Absil,, 2016; Lin et al.,, 2017; Zhao et al.,, 2016; Zeng and So,, 2017; Wang et al.,, 2023).
Limitations of existing RMC methods.
Despite extensive research on the RMC problem, current algorithms have various limitations: (i) Several RMC methods fail to recover the low rank matrix unless the number of observed entries is quite large; (ii) Some methods require as input the (often unknown) rank $r$ of the target matrix and fail when overparameterized even by just a single additional dimension (an input rank of $r+1$); (iii) Various methods fail to recover a matrix with a moderate condition number, as low as 5.
[…]
These shortcomings raise the following challenge: Develop a computationally efficient robust matrix completion method that can handle ill-conditioned matrices, overparameterization and requires only a small number of observed entries.
Our contributions.
In this work we make a step towards resolving this challenge. We propose a novel RMC method, denoted RGNMR, which overcomes the above limitations. Our proposed method is a simple iterative algorithm based on Gauss–Newton linearization and removal of entries suspected to be outliers. As seen in Figures 1, 2 and 3, empirically the performance of RGNMR is not affected by an overparameterized rank or by the condition number. Furthermore, RGNMR recovers underlying matrices with significantly fewer observed entries than existing methods. In Appendix B we show that RGNMR continues to outperform other methods under broader conditions, including both outliers and additive noise, non-uniform sampling and various other ranks, fraction of outliers, etc. Furthermore, we show that RGNMR perform well on a real data from computer vision, involving background extraction. In addition to its excellent empirical performance, in Section 3 we derive theoretical recovery guarantees for RGNMR. Specifically, we present two theorems. Theorem 3.1 states that given a suitable initialization, RGNMR recovers the target matrix at a linear rate. This theorem holds under the weakest known assumptions for factorization based RMC methods. Theorem 3.2 establishes that the initialization scheme we propose yields a suitable initial estimate.
2 The RGNMR algorithm
Algorithm 1 RGNMR
Input:
  • $\{X_{i,j}\mid(i,j)\in\Omega\}$ observed entries
  • $r$ rank of $L^{*}$
  • $k$ upper bound on the number of corrupted entries
  • $\begin{pmatrix}U_{0}\\ V_{0}\end{pmatrix}\in\mathbb{R}^{n_{1}+n_{2}\times r}$ initialization
  • $\Lambda_{0}$ initial estimate of the set of corrupted entries
  • $T$ maximal number of iterations
Output: $\hat{L}$ of rank $r$
for $t=0\ldots T-1$ do
$\begin{pmatrix}U_{t+1}\\ V_{t+1}\end{pmatrix}=\arg\min_{U,V}\|U_{t}V^{\top}+UV_{t}^{\top}-U_{t}V_{t}^{\top}-X\|_{F(\Omega\setminus\Lambda_{t})}^{2}$
$\Lambda_{t+1}=\arg\min_{\begin{subarray}{c}\Lambda\subset\Omega,|\Lambda|=k\end{subarray}}\|U_{t}V_{t+1}^{\top}+U_{t+1}V_{t}^{\top}-U_{t}V_{t}^{\top}-X\|_{F(\Omega\setminus\Lambda)}^{2}$
end for
return $P_{r}(U_{T-1}V_{T}^{\top}+U_{T}V_{T-1}^{\top}-U_{T-1}V_{T-1}^{\top})$
To describe RGNMR, we first introduce some notation. For a subset $\Omega\subseteq[n_{1}]\times[n_{2}]$, we denote by $\mathcal{P}_{\Omega}$ the following projection operator
$\mathcal{P}_{\Omega}(A)_{i,j}=\begin{cases}A_{i,j}&(i,j)\in\Omega,\\ 0&(i,j)\not\in\Omega.\\ \end{cases}$
For a matrix $A$ and a set $\Omega$, we denote $\|A\|^{2}_{F(\Omega)}=\|\mathcal{P}_{\Omega}(A)\|^{2}_{F}=\sum_{(i,j)\in\Omega}A_{i,j}^{2}$. Finally, we denote the set of corrupted entries by $\Lambda_{*}=\{(i,j)\in\Omega\,|\,S^{*}_{i,j}\not=0\}$ and the number of corrupted entries by $k^{*}=|\Lambda_{*}|$. Both $\Lambda_{*}$ and $k^{*}$ are unknown.
The optimization variables of RGNMR are two factor matrices $U\in\mathbb{R}^{n_{1}\times r},V\in\mathbb{R}^{n_{2}\times r}$ and a subset $\Lambda\subset\Omega$ which estimates the locations of the corrupted entries. RGNMR receives as input the following quantities; the subset of observed entries $\{X_{i,j}\mid(i,j)\in\Omega\}$; the rank of the target matrix $r$; an upper bound $k$ on the number of corrupted entries; initial values of factor matrices $(U_{0},V_{0})$, and of the set of corrupted entries $\Lambda_{0}\subset\Omega$.
At each iteration RGNMR performs the following two steps: (i) given the current set of suspected outlier entries $\Lambda$, update $U,V$ using the remaining entries $\mathcal{P}_{\Omega\setminus\Lambda}(X)$; (ii) given the updated matrix $L$, recompute the new set of suspected outliers $\Lambda$, by the $k$ entries with largest magnitude in $\mathcal{P}_{\Omega}(X-L)$.
Let us provide some motivation for the above two steps. If the estimated set of non corrupted entries, $\Omega\setminus\Lambda$, is large enough and contains only non-corrupted entries, namely $\Lambda_{*}\subset\Lambda$, then employing any vanilla matrix completion method would recover $L^{*}$ exactly from $\mathcal{P}_{\Omega\setminus\Lambda}(X)=\mathcal{P}_{\Omega\setminus\Lambda}(L^{*})$. Regarding the second step, consider an ideal case where at some iteration after step (i) RGNMR obtained $L=L^{*}$. Then at all the non-corrupted entries $(i,j)\in\Omega\setminus\Lambda_{*}$, the residual $X_{i,j}-L_{i,j}=0$ and $L^{*}$ is a fixed point of the algorithm. Therefore, given $L$ it is reasonable to estimate the corrupted entries by the entries with largest magnitude in $\mathcal{P}_{\Omega}(X-L)$. Iterating this process we hope to identify a sufficiently large subset of $\Omega\setminus\Lambda_{*}$ and consequently recover $L^{*}$.
We now describe the two steps of RGNMR. The first step builds upon the matrix completion approach of Zilber and Nadler, (2022). Formally, given the estimate $\Lambda_{t}$ of $\Lambda_{*}$ and the current factorization $(U_{t},V_{t})$, the updated matrices $\left(U_{t+1}V_{t+1}\right)$ are the solution of the following least squares problem,
$\begin{pmatrix}U_{t+1},V_{t+1}\end{pmatrix}=\arg\min_{U,V}\mathcal{L}^{t}_{{\Omega\setminus\Lambda_{t}}}(U,V)=\arg\min_{U,V}\|U_{t}V^{\top}+UV_{t}^{\top}-U_{t}V_{t}^{\top}-X\|_{F(\Omega\setminus\Lambda_{t})}.$ (1)
This problem is rank deficient with an infinite number of solutions. As in (Zilber and Nadler,, 2022), we choose the solution with minimal norm of the new estimate $\|U_{t+1}\|_{F}^{2}+\|V_{t+1}\|_{F}^{2}$.
For the second step, we construct a new estimate of $\Lambda_{*}$ by the $k$ largest residual entries in (1),
$\Lambda_{t+1}=\arg\min_{\begin{subarray}{c}\Lambda\subset\Omega,|\Lambda|=k\end{subarray}}\|U_{t}V_{t+1}^{\top}+U_{t+1}V_{t}^{\top}-U_{t}V_{t}^{\top}-X\|_{F(\Omega\setminus\Lambda)}^{2}.$ (2)
RGNMR iterates these two steps until convergence or until a maximal number of iterations $T$ is reached. Algorithm 1 presents an outline of RGNMR.
Remark 2.1.
The GNMR method of Zilber and Nadler, (2022) iterates Eq. (1) till convergence. If $\Omega\setminus\Lambda$ is large enough and contains only non corrupted entries, then after a few iterations GNMR would exactly recover $L^{*}$. However, since in the first few iterations the set $\Omega\setminus\Lambda$ often still includes a few corrupted entries, in our approach we re-estimate $\Lambda$ after each update of $(U_{t},V_{t})$ in (1). This yields considerable savings in terms of runtime.
2.1 Estimating the number of corrupted entries
One of the input parameters to RGNMR is the upper bound $k$ on the number of corrupted entries. At each iteration RGNMR removes the $k$ entries with largest residual error, and performs low-rank matrix completion on the remaining entries. For RGNMR to succeed, it is crucial for $k$ to be a tight upper bound. Indeed, if $k\gg k^{*}$ then too many entries are removed and on the remaining entries the corresponding matrix completion problem is ill posed.
As the true number of corrupted entries $k^{*}$ is unknown, it is important to develop a method to tightly upper bound it. Here we propose a scheme to do so. Our approach is motivated by the empirical observation that the estimates $\Lambda_{t}$ of RGNMR behave differently if $k\leq k^{*}$ or $k>k^{*}$. If $k\leq k^{*}$ then the estimates $\Lambda_{t}$ of the set of corrupted entries converge after several iterations. However, if $k>k^{*}$, the sets $\Lambda_{t}$ do not converge. This is illustrated in Figure 4 in the appendix, and we provide some intuition for this behavior below. Building upon this observation, we propose a binary search algorithm to estimate $k^{*}$, assuming it belongs to the range $[k_{\min},k_{\max}]$. In practice, we may take $k_{\min}=0$ and $k_{\max}=|\Omega|/2$ (namely at most 50% corrupted entries). At each step we run RGNMR with $k=\lfloor(k_{\min}+k_{\max})/2\rfloor$. If the sets $\Lambda_{t}$ converged, then we update $k_{\min}=k$. Otherwise, we set $k_{\max}=k$. After $\log_{2}(k_{\max}-k_{\min})$ steps we obtain a tight upper bound on $k^{*}$.
To provide intuition to why RGNMR behaves differently with $k\leq k^{*}$ or $k>k^{*}$, consider the following scenario. Assume that $k\leq k^{*}$ and that at some iteration, the set $\Lambda_{t}$ of $k$ entries with largest current residuals, satisfies that $\Lambda_{t}\subseteq\Lambda_{*}$. The next estimate $L_{t+1}$ of $L^{*}$ minimizes $\|\mathcal{P}_{\Omega\setminus\Lambda_{t}}(L-X)\|_{F}$, ignoring the entries in $\Lambda_{t}$, which by assumption are all outliers. Hence, $L_{t+1}$ tries to fit a low rank matrix to the remaining entries, and thus we expect the entries in $\Lambda_{t}$ to remain the $k$ largest residual entries. Therefore, $\Lambda_{t}=\Lambda_{t+1}$ and as the same argument holds for all subsequent iterations the sets $\Lambda_{t}$ converge. In contrast, if $k>k^{*}$ then at any iteration $t$, the set $\Lambda_{t}$ contains at least $k-k^{*}$ non-corrupted entries. Assume that at some iteration $t$, the method successfully detected all outliers, $\Lambda_{*}\subset\Lambda_{t}$. Then, if $k$ is not too large, the next estimate should be very close to the true low rank matrix $L_{t+1}\approx L^{*}$. As the computations are done in finite precision, the residuals at both $\Omega\setminus\Lambda_{t}$ as well as at $\Lambda_{t}\setminus\Lambda_{*}$ are not precisely zero but rather, due to rounding errors, appear as very small random values. Therefore, even though $\Lambda_{*}\subset\Lambda_{t+1}$ its remaining $k-k^{*}$ entries are not the same as those of $\Lambda_{t}\setminus\Lambda_{*}$, but rather change at each iteration $t$. Hence, the sets $\Lambda_{t}$ do not converge.
To demonstrate that our binary search scheme successfully bounds the number of corrupted entries we compare two variants of RGNMR. The first, denoted simply RGNMR, is given the true number of corrupted entries $k^{*}$ as input, $k=k^{*}$. The second variant, denoted RGNMR-BS, first upper bounds $k^{*}$ using our binary search scheme and then uses this upper bound $\hat{k}$ as input, $k=\hat{k}$. As illustrated in Figures 1, 2 and 3, RGNMR-BS performance is similar to that of RGNMR, which implies that our method obtains a sufficiently tight upper bound on $k^{*}$.
Remark 2.2.
Several RMC methods aim to remove the corrupted entries and require an estimate of $k^{*}$ (Yan et al.,, 2013; Yi et al.,, 2016). Since the intuition provided above holds also for these methods, a similar scheme to estimate $k^{*}$ can be incorporated there as well.
3 Recovery guarantees
Algorithm 2 RGNMR - modified
Input: All the inputs of RGNMR and the following additional parameters:
  • $\mu$ incoherence parameter of $L^{*}$
  • $\alpha$ a bound on the fraction of outliers in each row/column
  • $\gamma$ over-removal factor
  • $\delta$ neighborhood parameter
Output: $L^{*}$ of rank $r$
for $t=0\ldots T-1$ do
$\Lambda_{t}=\textnormal{support}\left(\mathcal{T}_{\gamma\alpha}\left(U_{t}V_{t}^{\top}-X,\Omega\right)\right)$
$\begin{pmatrix}U_{t+1}\\ V_{t+1}\end{pmatrix}=\arg\min_{U,V}\left\{\|U_{t}V^{\top}+UV_{t}^{\top}-U_{t}V_{t}^{\top}-X\|_{F(\Omega\setminus\Lambda_{t})}^{2}\middle|(U,V)\in\mathcal{B}_{\mu}\cap\mathcal{C}\big{(}U_{t},V_{t},\tfrac{\delta}{4^{t+1}}\big{)}\right\}$
end for
return $L_{T}=U_{T}V_{T}^{\top}$
Table 1: Recovery guarantees requirements of various RMC methods, up to multiplicative constant factors. The weakest conditions for each type of methods are in bold.
Method
Type
Method
Sample Complexity
($pn_{2}\geq$)
Corruption Rate
($\alpha\leq$)
Factorization Based Zheng and Lafferty, (2016) $\max\{{\mu r\log n_{1}},\mu^{2}r^{2}\kappa^{2}\}$ $\alpha=0,$ No Corruption
Tong et al., (2021) $p=1,$ Fully Observed $\frac{1}{r^{\frac{3}{2}}\mu\kappa}$
Cai et al., (2024) $p=1,$ Fully Observed $\frac{1}{r^{\frac{3}{2}}\mu\kappa}$
Yi et al., (2016) $\mu^{2}r^{2}\kappa^{4}\log n_{1}$ $\frac{1}{r\mu\kappa^{2}}$
Our Theorem 3.1 $\max\{\mu r\log n_{1},\mu^{2}r^{2}\kappa^{2}\}$ $\frac{1}{r\mu\kappa}$
Full Matrix Cherapanamjeri et al., (2017) $\mu^{2}r^{2}\log^{2}(\mu r\sigma_{1}^{*})\log^{2}n_{1}$ $\frac{1}{r\mu}$
Wang and Wei, (2024) $\mu^{3}r^{3}\kappa^{4}\log n_{1}$ $\frac{1}{r^{2}\mu^{2}\kappa^{2}}$
In this section we present recovery guarantees for RGNMR. Recall that the goal is to recover a rank $r$ matrix $L^{*}\in\mathbb{R}^{n_{1}\times n_{2}}$ from a subset $\Omega$ of the entries of $X=L^{*}+S^{*}$. We start with some notations and the assumptions made in our theoretical analysis.
Notations.
The $i$-th largest singular value of the matrix $L^{*}\in\mathbb{R}^{n_{1}\times n_{2}}$ is denoted by $\sigma_{i}^{*}$, and its condition number is denoted by $\kappa=\sigma^{*}_{1}/\sigma^{*}_{r}$. We denote the operator norm of a matrix (a.k.a. spectral norm) by $\|A\|_{op}$, its Frobenius norm by $\|A\|_{F}$, its $i$-th row by $A_{(i,\cdot)}$, its $j$ column by $A_{(\cdot,j)}$ its largest row norm by $\|A\|_{2,\infty}\equiv\max_{i}\|A_{(i,\cdot)}\|$ and its zero norm by $\|A\|_{0}=|\{(i,j)\in[n_{1}]\times[n_{2}]\mid A_{i,j}\not=0\}|$. We define the matrix $|A|$ by $|A|_{(i,j)}=|A_{(i,j)}|$. We denote by $P_{r}(A)$ the best rank $r$ approximation of $A$ in Frobenius norm, namely the projection to the subspace spanned by the $r$ singular vectors with largest singular values. We denote $\Omega_{(\cdot,j)}=\{i\mid(i,j)\in\Omega\}$, $\Omega_{(i,\cdot)}=\{j\mid(i,j)\in\Omega\}$ and $c_{j}=|\Omega_{(\cdot,j)}|$, $r_{i}=|\Omega_{(i,\cdot)}|$. The Procrustes distance between $Z_{1},Z_{2}\in\mathbb{R}^{n\times r}$ is
$d_{P}(Z_{1},Z_{2})$ $=\min\left\{\|Z_{1}-Z_{2}P\|_{F}\,\mid|\,P\in\mathbb{R}^{r\times r}\text{ is orthogonal}\right\}.$ (3)
For a vector $v$ we denote by $v^{(k)}$ its $k$-th largest entry in absolute value.
Assumptions.
For the RMC problem to be well-posed, we make three standard assumptions.
Assumption 1.
The underlying matrix $L^{*}$ is incoherent with incoherence parameter $\mu$. Formally, if $U\Sigma V^{\top}$ is the SVD of $L^{*}$ then
$\|U\|_{2,\infty}\leq\sqrt{\mu r/n_{1}},\quad\|V\|_{2,\infty}\leq\sqrt{\mu r/n_{2}}.$ (4)
For future use, we denote by $\mathcal{M}(n_{1},n_{2},r,\mu,\kappa)$ the set of $n_{1}\times n_{2}$ matrices of rank $r$, incoherence parameter $\mu$ and condition number $\kappa$.
Assumption 2.
[Bernoulli Model] Each entry in $X$ is independently observed with probability $p$. Hence, the number of observed entries is not fixed, but rather $|\Omega|\sim Bin(n_{1}n_{2},p)$.
Assumption 3.
In each row and column, the fraction of observed entries which are corrupted is bounded. Formally, for a known $10<\alpha<$ 1, we assume that $S^{*}\in\mathcal{S}_{\alpha}^{\Omega}$, where
$\mathcal{S}_{\alpha}^{\Omega}=$$\left\{S\in\mathbb{R}^{n_{1}\times n_{2}}\middle|\forall i:\|\mathcal{P}_{\Omega}(S)_{(i,\cdot)}\|_{0}\leq\alpha r_{i}\land\forall j:\|\mathcal{P}_{\Omega}(S)_{(\cdot,j)}\|_{0}\leq\alpha c_{j}\right\}.$ (5)
Assumption 1 was introduced by Candes and Recht, (2012). Assumptions 2 and 3 or variants thereof are common in theoretical analyses of matrix completion, see for example (Candès et al.,, 2011; Yi et al.,, 2016; Cherapanamjeri et al.,, 2017; Wang and Wei,, 2024).
Similar to other works on robust matrix completion, we derive guarantees for a slightly modified version of our proposed method. Specifically, we make the following two modifications. First, based on assumption 3 that the fraction of corrupted entries in each row and column is sufficiently small, we estimate the set of corrupted entries $\Lambda_{*}$ in a different way. To this end we introduce the following thresholding operator $\mathcal{T}_{\theta}(A,\Omega)$. Given an input matrix $A$, a subset $\Omega$, and a parameter $\theta\in(0,1)$, the operator $\mathcal{T}_{\theta}(A,\Omega)$ keeps only those entries $(i,j)\in\Omega$ which belong to the largest $\theta$-fraction of entries in both the respective row $i$ and column $j$, and zeros out all the remaining entries. In case of entries having identical magnitude, ties are broken arbitrarily. Formally,
$\mathcal{T}_{\theta}(A,\Omega)=\begin{cases}\mathcal{P}_{\Omega}(A)_{i,j}&|A_{i,j}|>\left[|\mathcal{P}_{\Omega}(A)|_{(i,\cdot)}\right]^{(\lceil\theta r_{i}\rceil)}\land|A_{i,j}|>\left[|\mathcal{P}_{\Omega}(A)|_{(\cdot,j)}\right]^{(\lceil\theta c_{j}\rceil)}\\ 0,&\text{otherwise}.\end{cases}$ (6)
At each iteration $t$, given the current estimate $L_{t}=U_{t}V_{t}^{\top}$, the estimated set of corrupted entries is
$\Lambda_{t}=\textnormal{support}\left(\mathcal{T}_{\theta}\left(L_{t}-X,\Omega\right)\right).$ (7)
For RGNMR to be able to remove all outliers, a necessary condition is that $\theta\geq\alpha$. In what follows, we choose $\theta=\gamma\alpha$ where $\gamma\geq 1$ is the over-removal factor.
The second modification to RGNMR is to constrain the update $(U_{t+1},V_{t+1})$ to be in the vicinity of the current pair of factor matrices $(U_{t},V_{t})$ and to have bounded row norms. Formally, for a parameter $\delta_{t}>0$ we define the $\delta_{t}$ neighborhood of the current estimate $(U_{t},V_{t})$ as
$\mathcal{C}{(U_{t},V_{t},\delta_{t})}=\left\{\begin{pmatrix}\tilde{U}\\ \tilde{V}\end{pmatrix}\in\mathbb{R}^{(n_{1}+n_{2})\times r}\,\middle|\,\|\tilde{U}-U_{t}\|^{2}_{F}+\|\tilde{V}-V_{t}\|^{2}_{F}\leq{\delta_{t}}\right\},$ (8)
and denote the subset of factor matrices with bounded row norms by
$\mathcal{B}_{\mu}$ =$=\left\{\begin{pmatrix}U\\ V\end{pmatrix}\in\mathbb{R}^{(n_{1}+n_{2})\times r}\,\middle|\,\|U\|_{2,\infty}\leq\sqrt{\frac{3\mu r\sigma_{1}^{*}}{n_{1}}},\quad\|V\|_{2,\infty}\leq\sqrt{\frac{3\mu r\sigma_{1}^{*}}{n_{2}}}\right\},$ (9)
where the constant $3$ is arbitrary. Instead of eq. 1, for a suitably chosen $\delta>0$, the modified RGNMR updates the factor matrices as follows,
$(U_{t+1},V_{t+1})=\arg\min\left\{\mathcal{L}^{t}_{\Omega\setminus\Lambda_{t}}(U,V)\middle|(U,V)\in\mathcal{B}_{\mu}\cap\mathcal{C}\left(U_{t},V_{t},\frac{\delta}{4^{t+1}}\right)\right\}.$ (10)
Similar constraints were employed by Zilber and Nadler, (2022) and Keshavan et al., (2010), in deriving theoretical guarantees for their (non-robust) matrix completion algorithms. As these constraints are quadratic, the above problem may be written as a regularized least squares problem with quadratic regularization terms. Hence, it can be solved computationally efficiently. The modified RGNMR is described in Algorithm 2.
Main theorems.
To state our recovery guarantees we introduce the following definitions. For a rank $r$ matrix $L^{*}$ with smallest singular value $\sigma^{*}_{r}$, we define the following sets, as in (Zilber and Nadler,, 2022). First, we denote all factorizations of rank-$r$ matrices with a bounded error from $L^{*}$ by
$\mathcal{B}_{\text{err}}(\epsilon)$ =$=\left\{\begin{pmatrix}U,V\end{pmatrix}\in\mathbb{R}^{(n_{1}+n_{2})\times r}\,\middle|\,\|UV^{\top}-L^{*}\|_{F}\leq\epsilon\sigma_{r}^{*}\right\}.$ (11)
In particular, we denote by $\mathcal{B}^{*}=\mathcal{B}_{\textnormal{err}}(0)$ the set of all exact factorizations of $L^{*}$,
=$\mathcal{B}^{*}=\left\{\begin{pmatrix}U,V\end{pmatrix}\in\mathbb{R}^{(n_{1}+n_{2})\times r}\,\middle|\,UV^{\top}=L^{*}\right\}.$ (12)
We say that $U,V$ are balanced if $U^{\top}U=V^{\top}V$, and measure the imbalance by $\|U^{\top}U-V^{\top}V\|_{F}$. We denote all the pairs of factor matrices which are approximately balanced by
$\mathcal{B}_{\textnormal{bln}}(\delta)$ =$=\left\{\begin{pmatrix}U,V\end{pmatrix}\in\mathbb{R}^{(n_{1}+n_{2})\times r}\,\middle|\,\|U^{\top}U-V^{\top}V\|_{F}\leq\delta\sigma_{r}^{*}\right\}.$ (13)
Our first theorem states that starting from a sufficiently accurate balanced initialization with bounded row norms, the modified RGNMR of Algorithm 2 recovers $L^{*}$ with high probability and with a linear convergence rate.
Theorem 3.1.
Let $X=L^{*}+S^{*}$, where $L^{*}\in\mathcal{M}(n_{1},n_{2},r,\mu,\kappa)$, and without loss of generality $n_{1}\geq n_{2}$. Let the set of observed entries $\Omega$ follow Assumption 2, and the corruption matrix $S^{*}$ satisfy Assumption 3 for some known $\alpha\in(0,1)$. For large enough absolute constants $C,c_{l},c_{e},c_{\alpha},c_{\gamma}$ the following holds: If the fraction of corrupted entries is small enough, $\alpha<\frac{1}{c_{\alpha}r\mu\kappa}$ and the probability to observe an entry is high enough, $p\geq\frac{C\mu r}{n_{2}}\max\{\log n_{1},\mu r\kappa^{2}\}$, then w.p. at least $1-\frac{6}{n_{1}}$, Algorithm 2 with parameter $\frac{25\sigma_{r}^{*}}{c_{e}^{2}\kappa}\leq\delta\leq\frac{\sigma_{r}^{*}}{c_{e}\kappa}$, over removal factor $c_{\gamma}\leq\gamma\leq\sqrt{c_{\alpha}}$ and an initialization $(U_{0},V_{0})\in\mathcal{B}_{\textnormal{err}}(\frac{1}{c_{e}\sqrt{\kappa}})\cap\mathcal{B}_{\textnormal{bln}}(\frac{1}{2c_{l}})\cap\mathcal{B}_{\mu}$ converges linearly to $L^{*}$. That is, after $t$ iterations the estimate $L_{t}=U_{t}V_{t}^{\top}$ satisfies
$\|L_{t}-L^{*}\|_{F}\leq\frac{1}{2^{t}}\frac{\sigma_{r}^{*}}{c_{e}\sqrt{\kappa}}.$ (14)
Our second theorem states that under suitable conditions it is possible to construct an accurate initialization that satisfies the requirements of Theorem 3.1. To do so we employ the spectral based initialization scheme proposed by Yi et al., (2016), followed by a normalization procedure on the rows of the factor matrices. Similar normalization procedures were employed by Zheng and Lafferty, (2016) and Zilber and Nadler, (2022) for their analysis of matrix completion without outliers. Our initialization scheme, outlined in Algorithm 3, is described in detail in Appendix A.
Theorem 3.2.
Let $X=L^{*}+S^{*}$, $L^{*}\in\mathcal{M}(n_{1},n_{2},r,\mu,\kappa)$ , $S^{*}\in\mathcal{S}_{\alpha}^{\Omega}$ and without loss of generality, $n_{1}\geq n_{2}$. Then for any $c_{l},c_{e}$ for large enough constants $c_{\alpha},C$ the following holds: If $\alpha\leq\frac{1}{c_{\alpha}\kappa^{2}r^{\frac{3}{2}}\mu}$ and $p\geq C\frac{\mu r^{2}\kappa^{4}\log n_{1}}{n_{2}}$ then w.p. at least $1-\frac{6}{n_{1}}$ Algorithm 3 outputs a pair $(U_{0},V_{0})$ that satisfies
$(U_{0},V_{0})\in\mathcal{B}_{\textnormal{err}}\left(\frac{1}{c_{e}\sqrt{\kappa}}\right)\cap\mathcal{B}_{\textnormal{bln}}\left(\frac{1}{2c_{l}}\right)\cap\mathcal{B}_{\mu}.$ (15)
Remark 3.1.
Theorem 3.1 holds for $\frac{25\sigma_{r}^{*}}{c_{e}^{2}\kappa}\leq\delta\leq\frac{\sigma_{r}^{*}}{c_{e}\kappa}$. In principle, to provide a valid value of $\delta$ in this interval requires knowledge of $\sigma_{1}^{*}$ and $\sigma_{r}^{*}$. We note that several other works also required knowledge of $\sigma_{1}^{*}$ or $\sigma_{r}^{*}$ or both for their methods, see (Yi et al.,, 2016; Sun and Luo,, 2016; Cherapanamjeri et al.,, 2017). In fact, for our approach it suffices to estimate these two singular values up to constant multiplicative factors. Moreover, within the proof of Theorem 3.2, we show that under its assumptions, it is indeed possible to estimate $\sigma_{1}^{*}$ and $\sigma_{r}^{*}$ up to constant factors.
Remark 3.2.
Theorems 3.1 and 3.2 require that the parameters $\mu$ and $\alpha$ are known. Similar assumptions were made in previous works, e.g. (Yi et al.,, 2016; Cherapanamjeri et al.,, 2017; Sun and Luo,, 2016; Zilber and Nadler,, 2022).
Comparison to other recovery guarantees.
We compare our theoretical results to those derived for both factorization based methods as well as methods that operate on the full matrix.
For the vanilla matrix completion problem, where the observed entries are not corrupted, the smallest sample complexity requirement for factorization based methods was derived by Zheng and Lafferty, (2016). Remarkably, Theorem 3.1 matches this result for RGNMR even in the presence of corrupted entries. The number of samples required by Theorem 3.1 is smaller than those required for other RMC methods (Yi et al.,, 2016; Wang and Wei,, 2024) and if $\kappa<C\log(\mu r\sigma_{1}^{*})\log n_{1}$, for a suitable constant $C$, it is also smaller than the requirements in (Cherapanamjeri et al.,, 2017).
In terms of the fraction $\alpha$ of corrupted entries, the recovery guarantee with the highest $\alpha$ for factorization based methods was derived by Tong et al., (2021) and by Cai et al., (2024), under the assumption that the matrix is fully observed. Theorem 3.1 improves upon this result by a factor of $\mathcal{O}(\sqrt{r})$, even though the matrix is only partially observed. Hence our recovery guarantee holds under the weakest known conditions for factorization based methods. For general RMC methods the recovery guarantee with the highest $\alpha$ was derived by Cherapanamjeri et al., (2017). Theorem 3.1 requires a smaller $\alpha$ than in (Cherapanamjeri et al.,, 2017) by a factor of $\mathcal{O}(\kappa)$. Table 1 summarizes the sample complexity and corruption rate requirements of different methods.
Next, for constructing a suitable initialization the requirements on $\alpha$ and $p$ in Theorem 3.2 are more stringent than those for the recovery guarantee of Theorem 3.1. This separation is common in other works on matrix completion (Zheng and Lafferty,, 2016; Yi et al.,, 2016; Zilber and Nadler,, 2022). In terms of sample complexity Theorem 3.2 still improves upon the results in (Yi et al.,, 2016; Wang and Wei,, 2024) and in some parameter regimes also upon the result in (Cherapanamjeri et al.,, 2017). In terms of the fraction of outliers the condition in Theorem 3.2 is more stringent than in (Yi et al.,, 2016) and (Cherapanamjeri et al.,, 2017). Since empirically RGNMR succeeds even with random initialization we believe the condition imposed by Theorem 3.1 can be relaxed. This would allow for less stringent conditions in Theorem 3.2. We leave this for future work.
Limitations and future research.
Although our theoretical analysis improves upon existing results, it does not fully account for the remarkable empirical performance of RGNMR. For example, the conditions in Theorem 3.1 on both $\alpha$ and $p$ depend on the condition number $\kappa$ while empirically the performance of RGNMR is not affected by $\kappa$. In addition, our analysis did not consider the case of overparameterization. These remains open directions for future theoretical analysis of RGNMR.
Another possible line of research is to employ RGNMR to the problem of robust matrix recovery (Li et al., 2020b, ; Ding et al.,, 2021; Ma and Fattahi,, 2023). As discussed in Remark 2.1, RGNMR is closely related to the GNMR method (Zilber and Nadler,, 2022). Since GNMR was proven to successfully handle the problem of vanilla matrix recovery, where no data is corrupted, we believe the same can be achieved for RGNMR in the presence of corrupted data.
4 Numerical results
This section details the simulations conducted to compare the various RMC methods. Their results are illustrated in Figures 1, 2, 3 and in Appendix B. Each instance consists of generating a random matrix $L^{*}\in\mathbb{R}^{n_{1}\times n_{2}}$ of a given rank $r$, condition number $\kappa$, subset of observed entries $\Omega\subseteq[n_{1}]\times[n_{2}]$ and a corruption matrix $S^{*}$. We generate $L^{*}$ following (Tong et al.,, 2021). Specifically, we construct $U\in\mathbb{R}^{n_{1}\times r},V\in\mathbb{R}^{n_{2}\times r}$ with entries $i.i.d.$ from the standard normal distribution and orthonormalize their columns. We then construct $\Sigma\in\mathbb{R}^{r\times r}$ as a diagonal matrix with $r$ evenly spaced values between $1$ and $\frac{1}{\kappa}$ on the diagonal. We set $L^{*}=U\Sigma V^{\top}$. Next, to generate $\Omega$ we follow (Zilber and Nadler,, 2022). Given an oversampling ratio $\rho$ we sample $\Omega$ of size $\lfloor\rho r(n_{1}+n_{2}-r)\rfloor$ randomly without replacement and verify that it contains at least $r$ entries in each column and row. To construct the corruption matrix $S^{*}$ we first sample its support $\Lambda_{*}\subseteq\Omega$ as follows. Given a fraction $\alpha$ of corrupted entries we sample $\lfloor\alpha\cdot|\Omega|\rfloor$ entries out of $\Omega$ randomly without replacement. We then verify that the set $\Omega\setminus\Lambda_{*}$ contains at least $r$ entries in each column and row. Similar to (Cai et al.,, 2024; Wang and Wei,, 2024), we sample the value of each corruption entry uniformly between $-\max_{i,j}|L^{*}_{i,j}|$ and $\max_{i,j}|L^{*}_{i,j}|$. Each method is then provided with the matrix $\mathcal{P}_{\Omega}(L^{*}+S^{*})$, the set of entries $\Omega$ and the matrix rank $r$.
We compare RGNMR to AOP (Yan et al.,, 2013), RMC (Cambier and Absil,, 2016), RPCA-GD (Yi et al.,, 2016), RRMC (Cherapanamjeri et al.,, 2017), HUB (Ruppel et al.,, 2020) and HOAT (Wang et al.,, 2023). In addition, as described in Section 2.1, we compare RGNMR to RGNMR-BS. Implementations of all methods except RRMC are publicly available. An implementation of RRMC was kindly provided to us by Wang and Wei, (2024). MATLAB and Python implementations of RGNMR and RGNMR-BS are given in the supplementary and will be made publicly available upon publication. All simulations were run on a 2.1GHz Intel Core i7 CPU with 32GB of memory.
Each algorithm was executed with its default parameters, with the following exceptions: (i) To improve RPCA-GD and HUB performance we increased their number of iterations by using stricter convergence threshold than their default; (ii) In RPCA-GD we employed an over removal factor $\gamma$ of $4$ as it significantly improved its performance in our simulations; (iii) We tuned the parameter $\xi$ in HOAT to $30$; (iv) We provided AOP the exact number of corrupted entries; (v) For RRMC we use the same parameters used in (Wang and Wei,, 2024).
In Appendix B we illustrate the performance of RGNMR under a broad set of additional scenarios. Specifically, Figure 5 shows that RGNMR can handle a large fraction of corrupted entries. Figure 6 shows that RGNMR outperforms other methods also under non-uniform sampling. Figure 7 presents its stability under additive noise and Figure 8 demonstrates that RGNMR can recover matrices with a high rank even if the number of observed entries is relatively small. Figure 9 illustrates that RGNMR performs well on a real dataset involving background extraction in a video. This is a standard benchmark for RMC methods (Yi et al.,, 2016; Cherapanamjeri et al.,, 2017; Huang et al.,, 2021; Cai et al.,, 2024). Finally, Figure 10 show the runtime of RGNMR as a function of matrix size. As the figure demonstrates, for an $n\times n$ matrix, the runtime of RGNMR grows quadratically with $n$.
Appendix A Initialization procedure
Algorithm 3 Initialization procedure for Robust Matrix Completion
Input:
  • $\{X_{i,j}\mid(i,j)\in\Omega\}$ observed entries
  • $\mu$ incoherence parameter of $L^{*}$
  • $r$ rank of $L^{*}$
  • $p$ probability of observing an entry
  • $\alpha$ the maximal fraction of outliers in each row/column
Output: $(U_{0},V_{0})$
$S_{init}=\mathcal{T}_{\alpha}{\left(\mathcal{P}_{\Omega}(X),\Omega\right)}$ // An estimate of $S^{*}$
set $Z=\begin{pmatrix}U\\ V\end{pmatrix}=\text{b-SVD}_{r}{\big{[}}\frac{1}{p}\left(\mathcal{P}_{\Omega}(X)-S_{init}\right){\big{]}}$
set $\eta_{1}=\sqrt{\frac{\mu r}{n_{1}}}\|Z\|_{op}$ and $\eta_{2}=\sqrt{\frac{\mu r}{n_{2}}}\|Z\|_{op}$
set $Z_{0}=\begin{pmatrix}R_{\eta_{1}}(U_{0})\\ R_{\eta_{2}}(V_{0})\end{pmatrix}$
return $(U_{0},V_{0})$
Theorem 3.1 shows that starting from a sufficiently accurate balanced initialization with bounded row norms RGNMR recovers $L^{*}$. Therefore, the goal of our initialization procedure is to construct such initialization. Our proposed scheme starts by constructing an initial estimator of $S^{*}$, which we denote by $S_{init}$. To this end, the function $\mathcal{T}_{\alpha}$ (defined in eq. 6) is applied to $\mathcal{P}_{\Omega}(X)$. Formally,
$S_{init}=\mathcal{T}_{\alpha}\left(\mathcal{P}_{\Omega}(X),\Omega\right).$ (16)
If the initial estimate of $S^{*}$ is accurate enough, $S_{init}\approx S^{*}$, then $\frac{1}{p}(\mathcal{P}_{\Omega}(X)-S_{init})\approx\mathbb{E}[\mathcal{P}_{\Omega}(L^{*})]$, should be a sufficiently accurate estimation of $L^{*}$. Denote by $\tilde{U}\tilde{\Sigma}\tilde{V}^{\top}$ the SVD of $P_{r}\left(\frac{1}{p}(\mathcal{P}_{\Omega}(X)-S_{init})\right)$. To obtain a balanced factorization of $\frac{1}{p}(\mathcal{P}_{\Omega}(X)-S_{init})$ we use the following operator,
[…]
Note that the output of this operator is a perfectly balanced factorization which satisfies $U^{\top}U=V^{\top}V$. To bound the row norms of the factor matrices we apply a clipping operator $R_{\eta}$ with a suitably chosen $\eta$. The operator is applied to a vector $\mathbf{x}$ and it is defined as
$R_{\eta}(\mathbf{x})=\begin{cases}\mathbf{x}&\|\mathbf{x}\|\leq\eta,\\ \eta\frac{\mathbf{x}}{\|\mathbf{x}\|}&\|\mathbf{x}\|>\eta.\end{cases}$ (18)
For a matrix $A$, we define $R_{\eta}(A)$ as the matrix obtained by applying $R_{\eta}$ to each of its rows. For every row $i$, $R_{\eta}(\mathbf{A})_{(i,\cdot)}=R_{\eta}(\mathbf{A}_{(i,\cdot)})$. Algorithm 3 outlines this initialization scheme.
Appendix B Additional simulations
Figure 5: Failure probability (±$\pm 1.96$SE) of RMC methods as a function of the fraction of corrupted entries $\alpha$. (left) Oversampling ratio of $\frac{|\Omega|}{r(n_{1}+n_{2}-r)}=8$; (right) Oversampling ratio of $\frac{|\Omega|}{r(n_{1}+n_{2}-r)}=12$. The underlying matrix $L^{*}$ if of size $3200\times 400$ has a rank $r=5$ and a condition number $\kappa=2$. Each point corresponds to 100 independent realizations.
Figure 6: Results of RMC methods as a function of the oversampling ratio under non uniform sampling. (left) Median rel-RMSE ; (right) Failure probability defined as $\mathbb{P}(\texttt{rel-RMSE}>10^{-3})$. The underlying matrix $L^{*}$ is of size $1000\times 1000$ has rank $r=5$ and condition number $\kappa=2$. The fraction of corrupted entries is $\alpha=5\%$. Each point corresponds to 50 independent realizations.
Figure 7: Performance of RMC methods in the presence of additive noise. The y-axis is the median rel-RMSE. The x-axis is the noise standard deviation. The underlying matrix $L^{*}$ is of size $3200\times 400$ has rank $r=5$ and condition number $\kappa=2$. The fraction of corrupted entries is $\alpha=5\%$. Each point corresponds to 50 independent realizations.
(a) RGNMR
(b) HUB
(c) RMC
(d) AOP
(e) RRMC
(f) HOAT
Figure 8: Performance of RMC methods as a function of the rank and the fraction of observed entries $\frac{|\Omega|}{n_{1}n_{2}}$. Deep blue corresponds to $0\%$ failure, white corresponds to $100\%$ failure. The matrix is of size $1000\times 1000$, the fraction of corrupted entries is $\alpha=10\%$ and the condition number is $\kappa=2$. Each point corresponds to 20 independent realizations.
(a) Original Image
(b) Sampled Image
(c) RMC
(d) AOP
(e) RGNMR
(f) RPCA-GD
(g) HOAT
Figure 9: Background extraction for “Hall” video data. The frames are recovered from $5\%$ of the original entries with an input rank of $r=1$.
Figure 10: Median run time of RGNMR and RGNMR -BS for a matrix of size $n\times n$ as a function of $n$. The matrix has rank $r=5$ and a condition number $\kappa=2$. The oversampling ratio is $\rho=\frac{|\Omega|}{r(2n-r)}=6$ and the fraction of corruption is $\alpha=5\%$. Each point corresponds to 20 independent realizations. Both $x$ and $y$ axes are on a log scale and the dashed lines have a slope of 2.
We present here results of additional simulations under various settings beyond those described in the main text.
Fraction of outliers.
Figure 5 shows the performance of various RMC methods as a function of the fraction of corrupted entries $\alpha$. We only compare methods that performed well at an oversampling ratio of $\frac{|\Omega|}{r(n_{1}+n_{2}-r)}=8$ with $\alpha=5\%$, see Figure 1, excluding RPCA-GD, HOAT and RRMC. As shown in the left figure if the number of observed entries is relatively small, at oversampling ratio of $8$, then RGNMR can handle a larger fraction of corrupted entries than other methods. In the right figure we show that if the number of observed entries is relatively large, oversampling ratio of $12$, then methods such as RMC and AOP can handle a larger number of corrupted entries than RGNMR. We note that if the condition number is high then RGNMR still outperforms these methods.
Non-uniform sampling.
In many applications, the entries of the matrix are not sampled uniformly (Meka et al.,, 2009). Hence we made simulations where the observed entries followed a power law sampling scheme similar to (Meka et al.,, 2009), described in the following lines. Given an oversampling ratio $\rho$ we define $w=\rho\cdot r\cdot(n_{1}+n_{2}-r)$. We construct two sequences $(\tilde{p}_{1}\dots\tilde{p}_{n_{1}})$ and $(\tilde{q}_{1}\dots\tilde{q}_{n_{2}})$ such that $\tilde{p}_{i}=i^{-\frac{2}{3}}$, $\tilde{q}_{i}=i^{-\frac{2}{3}}.$ We then normalize them to construct two new sequences
$p_{i}=w\cdot\frac{\tilde{p}_{i}}{\sum_{j}\tilde{p}_{j}},\quad q_{i}=w\cdot\frac{\tilde{q}_{i}}{\sum_{j}\tilde{q}_{j}}.$
Note that $\sum_{i}p_{i}=\sum_{j}q_{j}=w$. We sample each entry $(i,j)$ with probability $\frac{p_{i}q_{j}}{w}$. The expected number of observed entries is then $\mathbb{E}[\Omega]=w$. In Figure 6 we illustrate that RGNMR performs better than most methods when the entries are sampled under this scheme.
Outliers and additive noise.
In Figure 7 we illustrate RGNMR stability to additive noise. In this simulation, all observed entries are corrupted by additive white Gaussian noise with known standard deviation $\sigma$, in addition to the few outliers entries. In this scenario inliers entries are corrupted with random noise. Note that our method for upper bounding $k^{*}$ is based on the assumption that the error in those entries is the result of rounding errors, see 2.1. To overcome this problem, when searching for an upper bound on $k^{*}$ we terminated RGNMR once $\frac{\|L_{t}-X\|_{F(\Omega\setminus\Lambda_{t})}}{\|X\|_{F(\Omega\setminus\Lambda_{t})}}\leq\sqrt{\sigma}$.
High rank matrices.
In Figure 8 we compare the performance of various RMC method as a function of the rank of the target matrix $L^{*}$. Following (Huang et al.,, 2021; Wang and Wei,, 2024) we fix the corruption rate $\alpha=0.1$ and vary the rank and the fraction of observed entries $\frac{|\Omega|}{n_{1}n_{2}}$. As shown, RGNMR successfully recovers $L^{*}$ from a small fraction of the observed entries even if the rank of $L^{*}$ is relatively high.
Real data: Background extraction from video.
In Figure 9 we illustrate RGNMR performance for video background extraction. We use the data from (Li et al.,, 2004), kindly provided to us by the authors of (Huang et al.,, 2021). The data contains a sequence of grayscale frames. By stacking the columns of each frame of the video into a long vector, we obtain a matrix whose columns correspond to the frames. This matrix can be decomposed to a low rank matrix corresponding to the static background plus a sparse matrix corresponding to the moving foreground. Following (Yi et al.,, 2016; Cherapanamjeri et al.,, 2017; Huang et al.,, 2021; Cai et al.,, 2024) we sample uniformly at random $5\%$ of the matrix entries. All method are then given the sampled entries and an input rank of $r=1$. As shown RGNMR performs well on real data and successfully extracts the background.
Runtime.
Finally, In Figure 10 we illustrate that RGNMR scales well with the matrix size. We show that for a matrix of size $n\times n$ the runtime of RGNMR and RGNMR-BS increases quadratically with $n$.
Appendix C Proofs of Theorems 3.1 and 3.2
To prove Theorems 3.1 and 3.2, we make use of several auxiliary lemmas. These are outlined in subsections C.1 and C.3. The proofs of the two theorems appear in subsections C.2 and C.4.
C.1 Auxiliary Lemmas For Theorem 3.1
The first three lemmas are results from (Zilber and Nadler,, 2022). The first one is a combination of Lemmas SM5.3 and SM5.5 from (Zilber and Nadler,, 2022).
Lemma C.1.
Let $L^{*}\in\mathcal{M}(n_{1},n_{2},r,\mu,\kappa)$ and let $\Omega$ follow Assumption 2 with probability $p$. Let $\epsilon\in(0,1)$. There exist constants $C,c_{l},c_{e}$ such that the following holds: If $p\geq C\max\{\frac{\log n_{1}}{n_{2}},\frac{\mu^{2}r^{2}\kappa^{2}}{n_{2}\epsilon^{4}}\}$. Then w.p. at least $1-\frac{2}{n_{1}^{5}}$, for any $(U,V)\in\mathcal{B}_{\textnormal{err}}({\epsilon}/{c_{e}})\cap\mathcal{B}_{\textnormal{bln}}({1}/{c_{l}})\cap\mathcal{B}_{\mu}$ with $L=UV^{\top}$ there exists $(U^{*},V^{*})\in\mathcal{B}^{*}\cap\mathcal{B}_{\mu}$ such that
$\|U-U^{*}\|_{F}^{2}+\|V-V^{*}\|_{F}^{2}$ $\leq\frac{25}{4\sigma_{r}^{*}}\|UV^{\top}-L^{*}\|_{F}^{2},$ (19a)
$\frac{1}{\sqrt{p}}\|(U-U^{*})(V-V^{*})^{\top}\|_{F(\Omega)}$ $\leq\frac{\epsilon}{6}\|L-L^{*}\|_{F}.$ (19b)
The second lemma is Lemma SM5.4 in (Zilber and Nadler,, 2022).
Lemma C.2 (uniform RIP for matrix completion).
Let $L^{*}\in\mathcal{M}(n_{1},n_{2},r,\mu,\kappa)$ and let $\Omega$ follow Assumption 2 with probability $p$. Let $\epsilon\in(0,1)$. There exist constants $C,c_{l},c_{e}$ such that the following holds: If $p\geq C\frac{\log n_{1}\mu r}{n_{2}\epsilon^{2}}$. Then w.p. at least $1-\frac{3}{n_{1}^{3}}$, for all matrices $L=UV^{\top}$ where $(U,V)\in\mathcal{B}_{\textnormal{err}}(\epsilon/c_{e})\cap\mathcal{B}_{\textnormal{bln}}(1/c_{l})\cap\mathcal{B}_{\mu}$, the following holds,
$(1-\epsilon)\|L-L^{*}\|_{F}\leq\tfrac{1}{\sqrt{p}}\|\mathcal{P}_{\Omega}(L-L^{*})\|\leq(1+\epsilon)\|L-L^{*}\|_{F}.$ (20)
The third lemma is Lemma SM5.6 in (Zilber and Nadler,, 2022), which in turn is a direct consequence of Lemma 7.1 in (Keshavan et al.,, 2010). In (Zilber and Nadler,, 2022) the result is stated with an unspecified constant $c$. Tracking its proof it can be shown that it holds with $c=2$, which is how the lemma is stated below.
Lemma C.3.
Let $\Omega$ follow Assumption 2 with probability $p$. There exist a constant $C$ such that the following holds for any $\mu,t,\epsilon>0$: If $p\geq C\max\{\frac{\log n_{1}}{n_{2}},\frac{\mu^{2}r^{2}}{n_{2}\epsilon^{4}}\}$. Then w.p. at least $1-\frac{2}{n_{1}^{5}}$, for any $(U,V)\in\mathbb{R}^{(n_{1}+n_{2})\times r}$ such that
$\|U\|_{2,\infty}\leq 4\sqrt{\mu rt/n_{1}},\quad\|V\|_{2,\infty}\leq 4\sqrt{\mu rt/n_{2}},$ (21)
we have
$\frac{1}{p}\|UV^{\top}\|^{2}_{F(\Omega)}\leq\frac{\|U\|_{F}^{2}+\|V\|_{F}^{2}}{2}\left(2\left(\|U\|_{F}^{2}+\|V\|_{F}^{2}\right)+t\epsilon^{2}\right).$ (22)
The following two lemmas provide bounds on the estimates constructed by RGNMR at each iteration. The first lemma states that starting from an approximately balanced and sufficiently accurate estimate of $L^{*}$, if each new pair of factor matrices is sufficiently close to the previous one, then all factor matrices continue to be approximately balanced, and remain not too far from $L^{*}$. Its proof appears in Appendix D.1.
Lemma C.4.
Let $L^{*}\in\mathcal{M}(n_{1},n_{2},r,\mu,\kappa)$ with $r$-th singular value $\sigma_{r}^{*}$. For large enough constants $c_{l},c_{e}$ the following holds: If $(U_{0},V_{0})\in\mathcal{B}_{\textnormal{err}}(\frac{1}{c_{e}\sqrt{\kappa}})\cap\mathcal{B}_{\textnormal{bln}}(\frac{1}{2c_{l}})\cap\mathcal{B}_{\mu}$ and for every $11\leq s\leq t+$ 1 , $(U_{s},V_{s})\in\mathcal{C}(U_{s-1},V_{s-1},\frac{\sigma_{r}^{*}}{4^{s}c_{e}\kappa})$ , then $(U_{t+1},V_{t+1})\in\mathcal{B}_{\textnormal{err}}(\frac{13}{\sqrt{c_{e}}})\cap\mathcal{B}_{\textnormal{bln}}(\frac{1}{c_{l}})$.
The second lemma bounds the distance between $L^{*}$ and the updated estimate $L_{t+1}$, as a function of various quantities of the current estimate $L_{t}$. Its proof appears in Appendix D.2.
Lemma C.5.
Let $L_{t}=U_{t}V_{t}^{\top}$ be the estimate of $L^{*}$ at iteration $t$. Let $L_{t+1}=U_{t+1}V_{t+1}^{\top}$ be the updated estimate, where $(U_{t+1},V_{t+1})$ are computed by (10) with some $\delta>0$. Assume that $L_{t}$ is sufficiently close to $L^{*}$, so that the set $\mathcal{A}_{t}=\mathcal{B}^{*}\cap\mathcal{B}_{\mu}\cap\mathcal{C}\left(U_{t},V_{t},\frac{\delta}{4^{t+1}}\right)$ is non empty. Let $(U^{*},V^{*})\in\mathcal{A}_{t}$ and denote $\Delta U_{t}^{*}=U_{t}-U^{*}$, $\Delta U_{t+1}=U_{t+1}-U_{t}$ with similar definitions for $\Delta V_{t}^{*}$ and $\Delta V_{t+1}$. Then, the error on the non removed entries $\Omega_{t}=\Omega\setminus\Lambda_{t}$ of the updated estimate satisfies
$\begin{split}\|L_{t+1}-L^{*}\|_{F({\Omega_{t})}}\leq&\sqrt{2}\|\Delta U_{t}^{*}\Delta V_{t}^{*\top}\|_{F({\Omega_{t}})}+\sqrt{2}\|\Delta U_{t+1}\Delta V_{t+1}^{\top}\|_{F({\Omega_{t})}}\\ &+(1+\sqrt{2})\|S^{*}\|_{F(\Omega_{t}\cap\Lambda_{*})}.\end{split}$ (23)
Before we present the last two lemmas, whose proofs appears in Appendix D.3, recall that at each iteration $t$, RGNMR removes from $\Omega$ the entries in $\Lambda_{t}=\text{support}(\mathcal{T}_{\gamma\alpha}\left(L_{t}-X,\Omega\right))$. If RGNMR was unable to identify all the corrupted entries, then the set $(\Omega\setminus\Lambda_{t})\cap\Lambda_{*}$, where $\Lambda_{*}=\mbox{support}(S^{*})$, is not empty. The following lemma bounds the magnitude of the entries of $S^{*}$ in $(\Omega\setminus\Lambda_{t})\cap\Lambda_{*}$, for matrices $L_{t}$ that satisfy a suitable condition.
Lemma C.6.
Let $X=L^{*}+S^{*}$ where $L^{*}\in\mathcal{M}(n_{1},n_{2},r,\mu,\kappa)$ and the corruption matrix $S^{*}$ satisfies Assumption 3 for some known $10<\alpha<$ 1. Let the set of observed entries $\Omega$ follow Assumption 2 with $p\geq C\frac{\log n_{1}}{n_{2}}$ for some constant $C$. Let $L=UV^{\top}$ be an estimate of $L^{*}$ such that $(U,V)\in\mathcal{B}_{\mu}$. Suppose that the set of corrupted entries is estimated by $\Lambda=\text{support}\left(\mathcal{T}_{\gamma\alpha}\left(L-X,\Omega\right)\right)$ with an over removal factor of $1<\gamma\leq\frac{1}{\alpha}$. If there exists a factorization $(U^{*},V^{*})\in\mathcal{B}^{*}\cap\mathcal{B}_{\mu}$ of $L^{*}$ such that
$\|U-U^{*}\|_{F}^{2}+\|V-V^{*}\|_{F}^{2}$ $\leq\frac{25}{4\sigma_{r}^{*}}\|L-L^{*}\|_{F}^{2},$ (24)
then the magnitude of the remaining corrupted entries in $(\Omega\setminus\Lambda)\cap\Lambda^{*}$, is bounded as follows,
$\|S^{*}\|_{F((\Omega\setminus\Lambda)\cap\Lambda_{*})}\leq 18\sqrt{p\alpha\mu r\kappa}\lVert L-L^{*}\rVert_{F}+\sqrt{\frac{2}{\gamma-1}}\lVert L-L^{*}\rVert_{F(\Omega)}.$ (25)
The last lemma bounds the error term $\|L-L^{*}\|_{F{(\Lambda_{t})}}$ for a matrix $L$ that satisfies a suitable condition.
Lemma C.7.
Let $L^{*}\in\mathcal{M}(n_{1},n_{2},r,\mu,\kappa)$ and let the set of observed entries $\Omega$ follow Assumption 2 with $p\geq C\frac{\log n_{1}}{n_{2}}$ for some constant $C$. Let $L=UV^{\top}$ be an estimate of $L^{*}$ such that $(U,V)\in\mathcal{B}_{\mu}$ and assume there exists a factorization $(U^{*},V^{*})\in\mathcal{B}^{*}\cap\mathcal{B}_{\mu}$ of $L^{*}$. If $\Lambda=\text{support}\left(\mathcal{T}_{\gamma\alpha}\left(A,\Omega\right)\right)$, for some matrix $A$, then
$\|L-L^{*}\|_{F(\Lambda)}\leq 27\sqrt{3}\gamma\alpha p\mu r\sigma_{1}^{*}(\lVert U-U^{*}\rVert_{F}^{2}+\lVert V-V^{*}\rVert_{F}^{2}).$ (26)
C.2 Proof of Theorem 3.1
The proof relies on lemmas C.1-C.7 above. We first note that since $\Omega$ follows Assumption 2 with $p\geq\frac{C\mu r}{n_{2}}\max\{\log n_{1},\mu r\kappa^{2}\}$, we may apply these lemmas. Specifically, for a large enough $C$ the conditions required by C.1 and C.2 with $\epsilon\geq\frac{1}{20}$ hold and the condition of Lemma C.3 holds with $\epsilon\geq\frac{1}{8\sqrt{c_{e}\kappa}}$.
Proof of Theorem 3.1.
We prove by induction on the iteration step $t$ that
$(U_{t},V_{t})\in\mathcal{B}_{\textnormal{err}}\left(\frac{1}{2^{t}c_{e}\sqrt{\kappa}}\right)\cap\mathcal{B}_{\textnormal{bln}}\left(\frac{1}{c_{l}}\right)\cap\mathcal{B}_{\mu}.$ (27)
Eq. eq. 14 then follows immediately by the definition of $\mathcal{B}_{\textnormal{err}}$, see eq. 11.
Eq.(27) for $t=0$ follows from the assumption that the initialization is sufficiently accurate, balanced and has bounded row norms $(U_{0},V_{0})\in\mathcal{B}_{\textnormal{err}}(\frac{1}{c_{e}\sqrt{\kappa}})\cap\mathcal{B}_{\textnormal{bln}}(\frac{1}{2c_{l}})\cap\mathcal{B}_{\mu}$. Next we assume $(U_{t},V_{t})$ satisfies eq. 27 and prove that the updated matrices also satisfy this equation. Since at all intermediate steps $11\leq s\leq t+$ 1, $(U_{s},V_{s})$ are updated by eq. 10 , $(U_{s},V_{s})\in\mathcal{C}(U_{s-1},V_{s-1},\frac{\sigma_{r}^{*}}{4^{s}c_{e}\kappa})\cap\mathcal{B}_{\mu}$. Hence by Lemma C.4, $(U_{t+1},V_{t+1})\in\mathcal{B}_{\textnormal{bln}}(\frac{1}{c_{l}})\cap\mathcal{B}_{\mu}$.
It remains to prove that $(U_{t+1},V_{t+1})\in\mathcal{B}_{\textnormal{err}}(\frac{1}{2^{t+1}c_{e}\sqrt{\kappa}})$, namely that $\|L_{t+1}-L^{*}\|_{F}\leq\frac{\sigma_{r}^{*}}{2^{t+1}c_{e}\sqrt{\kappa}}$. The proof consist of two parts. First we bound the error restricted to the set $\Omega_{t}=\Omega\setminus\Lambda_{t}$, i.e. $\|L_{t+1}-L^{*}\|_{F({\Omega_{t})}}$. In the second part we use this result to bound the error on the entire matrix.
For the first part, to bound $\|L_{t+1}-L^{*}\|_{F({\Omega_{t})}}$ we apply Lemma C.5. To this end we first prove its conditions hold. In particular that the set $\mathcal{A}_{t}=\mathcal{B}^{*}\cap\mathcal{B}_{\mu}\cap\mathcal{C}\left(U_{t},V_{t},\frac{\delta}{4^{t+1}}\right)$ is non empty. Indeed this follows by applying Lemma C.1. Specifically, since $(U_{t},V_{t})\in\mathcal{B}_{\textnormal{err}}\left(\frac{1}{2^{t}c_{e}\sqrt{\kappa}}\right)\cap\mathcal{B}_{\textnormal{bln}}\left(\frac{1}{c_{l}}\right)\cap\mathcal{B}_{\mu}$, for a large enough $c_{e}$, $\left(U_{t},V_{t}\right)$ satisfy the conditions of Lemma C.1 with $\epsilon=\frac{1}{20}$. Consequently, w.p. at least $1-\frac{2}{n_{1}^{5}}$, there exists a pair $(U^{*},V^{*})\in\mathcal{B}^{*}\cap\mathcal{B}_{\mu}$ that satisfy eqs. 19a and 19b with $U=U_{t},V=V_{t}$. Combining eq. 19a with the assumptions $\frac{25\sigma_{r}^{*}}{c_{e}^{2}\kappa}\leq\delta$ and $\|L_{t}-L^{*}\|\leq\frac{\sigma_{r}^{*}}{2^{t}c_{e}\sqrt{\kappa}}$ yields
$\|U_{t}-U^{*}\|_{F}^{2}+\|V_{t}-V^{*}\|_{F}^{2}$ $\leq\frac{25}{4\sigma_{r}^{*}}\|L_{t}-L^{*}\|_{F}^{2}\leq\frac{25}{4\sigma_{r}^{*}}\frac{\sigma_{r}^{*2}}{4^{t}c_{e}^{2}\kappa}\leq\frac{\delta}{4^{t+1}}.$
Therefore $(U^{*},V^{*})\in\mathcal{B}^{*}\cap\mathcal{B}_{\mu}\cap\mathcal{C}\left(U_{t},V_{t},\frac{\delta}{4^{t+1}}\right)$.
We can now apply Lemma C.5 which gives
$\begin{split}\|L_{t+1}-L^{*}\|_{F({\Omega_{t})}}\leq&\sqrt{2}\left(\underbrace{\|\Delta U_{t}^{*}\Delta V_{t}^{*}\|_{F({\Omega_{t}})}}_{T_{1}}+\underbrace{\|\Delta U_{t+1}\Delta V_{t+1}^{\top}\|_{F({\Omega_{t})}}}_{T_{2}}\right)\\ &+(1+\sqrt{2})\underbrace{\|S^{*}\|_{F(\Omega_{t}\cap\Lambda_{*})}}_{T_{3}}.\end{split}$ (28)
We upper bound the term $T_{1},T_{2}$ and $T_{3}$. The first term $T_{1}$ is bounded by eq. 19b of Lemma C.1 with $\epsilon=\frac{1}{20}$.
$T_{1}=\|\Delta U_{t}^{*}\Delta V_{t}^{*}\|_{F({\Omega_{t}})}\leq\|\Delta U_{t}^{*}\Delta V_{t}^{*}\|_{F(\Omega)}\leq\frac{\sqrt{p}}{120}\|L_{t}-L\|_{F}\leq\frac{\sqrt{p}\sigma_{r}^{*}}{2^{t+6}c_{e}\sqrt{\kappa}}.$ (29)
To bound $T_{2}$ we note that
$\left(\frac{1}{\sqrt{p}}T_{2}\right)^{2}$ $=\frac{1}{p}\|\Delta U_{t+1}\Delta V_{t+1}^{\top}\|^{2}_{F({\Omega_{t})}}\leq\frac{1}{p}\|\Delta U_{t+1}\Delta V_{t+1}^{\top}\|^{2}_{F({\Omega)}}$
To bound the above term we apply Lemma C.3. We first show that $(\Delta U_{t+1},\Delta V_{t+1})$ satisfies LABEL:{eq:KMO10_lemma71_consequence_assumption}. Since $(U_{t+1},V_{t+1})$ and $(U_{t},V_{t})$ are in $\mathcal{B}_{\mu}$ then
$\|\Delta U_{t+1}\|_{2,\infty}\leq\|U_{t+1}\|_{2,\infty}+\|U_{t}\|_{2,\infty}\leq 2\sqrt{\frac{3\mu r\sigma_{r}^{*}}{n_{1}}}\leq 4\sqrt{\frac{\mu r\sigma_{r}^{*}}{n_{1}}},$
and a similar inequality holds for $\Delta V_{t+1}$. Hence LABEL:{eq:KMO10_lemma71_consequence_assumption} holds with $t=\sigma_{r}^{*}$. Consequently by eq. 22 of C.3 with $\epsilon=\frac{1}{8\sqrt{c_{e}\kappa}}$, w.p. at least $1-\frac{2}{n_{1}^{5}}$
$\frac{1}{p}\|\Delta U_{t+1}\Delta V_{t+1}^{\top}\|^{2}_{F({\Omega)}}\leq\frac{\|\Delta U_{t+1}\|^{2}_{F}+\|\Delta V_{t+1}\|^{2}_{F}}{2}\left[2\left(\|\Delta U_{t+1}\|^{2}_{F}+\|\Delta V_{t+1}\|^{2}_{F}\right)+\frac{\sigma_{r}^{*}}{64c_{e}\kappa}\right].$
Next we explicitly bound $\|\Delta U_{t+1}\|^{2}_{F}+\|\Delta V_{t+1}\|^{2}_{F}$. Since $(U_{t+1},V_{t+1})\in\mathcal{C}\left(U_{t},V_{t},\frac{\delta}{4^{t+1}}\right)$, $\|\Delta U_{t+1}\|^{2}_{F}+\|\Delta V_{t+1}\|^{2}_{F}\leq\frac{\delta}{4^{t+1}}$. Hence
$T_{2}$ $\leq\sqrt{p}\sqrt{\frac{\delta}{2\cdot 4^{t+1}}\left[2\frac{\delta}{4^{t+1}}+\frac{\sigma_{r}^{*}}{64c_{e}\kappa}\right]}\leq\sqrt{p}\sqrt{\frac{\delta^{2}}{2\cdot 4^{2t+2}}+\frac{\delta\sigma_{r}^{*}}{128c_{e}\cdot 4^{t+1}\kappa}}$
Since $\delta\leq\frac{\sigma_{r}^{*}}{c_{e}\kappa}$
$T_{2}$ $\leq\sqrt{p}\sqrt{\frac{\sigma_{r}^{*2}}{c_{e}^{2}\kappa^{2}\cdot 2^{4t+4}}+\frac{\sigma_{r}^{*2}}{c_{e}^{2}\kappa^{2}\cdot 2^{2t+9}}}\leq\sqrt{p}\sqrt{\frac{\sigma_{r}^{*2}}{c_{e}^{2}\kappa^{2}\cdot 2^{2t+3.8}}}\leq\frac{\sqrt{p}\sigma_{r}^{*}}{c_{e}\sqrt{\kappa}\cdot 2^{t+1.9}}.$ (30)
In order to bound $T_{3}$ we invoke Lemma C.6. The condition of this Lemma, eq. 24, holds since we proved that $(U_{t},V_{t})$ satisfy eq. 19a of Lemma C.1. Since $\Lambda_{t}=\textnormal{support}\left(\mathcal{T}_{\theta}\left(L_{t}-X,\Omega\right)\right)$ we can employ Lemma C.6 which gives
$T_{3}\leq 18\sqrt{p\alpha\mu r\kappa}\lVert L_{t}-L^{*}\rVert_{F}+\sqrt{\frac{2}{\gamma-1}}\lVert L_{t}-L^{*}\rVert_{F(\Omega)}.$ (31)
Since $(U_{t},V_{t})\in\mathcal{B}_{\textnormal{err}}(\frac{1}{2^{t}c_{e}\sqrt{\kappa}})$, for a large enough $c_{e}$, $(U_{t},V_{t})$ satisfy the conditions of Lemma C.2 with $\epsilon=\frac{1}{20}$. Therefore by eq. 20 w.p. at least $1-\frac{3}{n_{1}^{3}}$,
$\sqrt{\frac{2}{\gamma-1}}\lVert L_{t}-L^{*}\rVert_{F(\Omega)}\leq(1+\frac{1}{20})\sqrt{\frac{2p}{\gamma-1}}\lVert L_{t}-L^{*}\rVert_{F}\leq\sqrt{\frac{3p}{\gamma-1}}\lVert L_{t}-L^{*}\rVert_{F}.$ (32)
Combining eq. 31 and eq. 32 gives,
$T_{3}\leq\sqrt{p}\left(18\sqrt{\alpha\mu r\kappa}+\sqrt{\frac{3}{\gamma-1}}\right)\lVert L_{t}-L^{*}\rVert_{F}.$
Since $\alpha<\frac{1}{c_{\alpha}\mu r\kappa},c_{\gamma}\leq\gamma$ and by the induction hypothesis $\lVert L_{t}-L^{*}\rVert_{F}\leq\frac{\sigma_{r}^{*}}{2^{t}c_{e}\sqrt{\kappa}}$, then
$T_{3}\leq\sqrt{p}\left(18\sqrt{\alpha\mu r\kappa}+\sqrt{\frac{3}{\gamma-1}}\right)\frac{\sigma_{r}^{*}}{2^{t}c_{e}\sqrt{\kappa}}\leq\sqrt{p}\left(\frac{18}{\sqrt{c_{\alpha}}}+\sqrt{\frac{3}{c_{\gamma}-1}}\right)\frac{\sigma_{r}^{*}}{2^{t}c_{e}\sqrt{\kappa}}.$
For large enough $c_{\alpha}$ and $c_{\gamma}$, this yields
$T_{3}\leq\frac{\sqrt{p}\sigma_{r}^{*}}{2^{t+5}c_{e}\sqrt{\kappa}(1+\sqrt{2})}.$ (33)
Inserting the bounds on $T_{1},T_{2}$ and $T_{3}$ from equations (29), (30) and (33) into eq. 28 gives
$\|L_{t+1}-L^{*}\|_{F({\Omega_{t})}}\leq\frac{\sqrt{p}\sigma_{r}^{*}}{2^{t+1}c_{e}\sqrt{\kappa}}\left(\frac{1}{2^{4.5}}+\frac{1}{2^{0.4}}+\frac{1}{2^{4}}\right)\leq\frac{9}{10}\frac{\sqrt{p}\sigma_{r}^{*}}{2^{t+1}c_{e}\sqrt{\kappa}}.$ (34)
Eq. (34) provides a bound on $L_{t+1}-L^{*}$, but only on the set $\Omega_{t}$. In what follows we show that this is sufficient for bounding the overall error on all entries, $\|L_{t+1}-L\|_{F}$. Since $\Omega_{t}=\Omega\setminus\Lambda_{t}$, then
$\lVert L_{t+1}-L^{*}\rVert^{2}_{F(\Omega_{t})}=\lVert L_{t+1}-L^{*}\rVert^{2}_{F(\Omega)}-\lVert L_{t+1}-L^{*}\rVert^{2}_{F(\Lambda_{t})}.$ (35)
To upper bound $\lVert L_{t+1}-L^{*}\rVert^{2}_{F(\Lambda_{t})}$ we apply Lemma C.7, with $(U_{t+1},V_{t+1})$ which indeed by their definition in eq. 10, belong to $B_{\mu}$, and with a specific factorization $(U^{*},V^{*})$ specified below. Specifically, we show that there exists a pair $(U^{*},V^{*})\in\mathcal{B}^{*}\cap\mathcal{B}_{\mu}$ that is in the vicinity of $\left(U_{t+1},V_{t+1}\right)$. By Lemma C.4 $\left(U_{t+1},V_{t+1}\right)\in\mathcal{B}_{\textnormal{err}}(\frac{13}{\sqrt{c_{e}}})\cap\mathcal{B}_{\textnormal{bln}}(\frac{1}{c_{l}})\cap\mathcal{B}_{\mu}$. Therefore, for a large enough $c_{e}$ Lemma C.1 guarantees the existence of a factorization $(U^{*},V^{*})\in\mathcal{B}^{*}\cap\mathcal{B}_{\mu}$ that also satisfy eq. 19a with $L=L_{t+1}=U_{t+1}V_{t+1}^{\top}$. Hence, by eq. 26 of Lemma C.7
$\lVert L_{t+1}-L^{*}\rVert^{2}_{F(\Lambda_{t})}\leq 27\sqrt{3}\gamma\alpha p\mu r\sigma_{1}^{*}(\lVert U_{t+1}-U^{*}\rVert_{F}^{2}+\lVert V_{t+1}-V^{*}\rVert_{F}^{2}).$ (36)
Combining eq. 19a and the assumptions $\alpha\leq\frac{1}{c_{\alpha}r\mu\kappa},\gamma\leq\sqrt{c_{\alpha}}$ gives
$\lVert L_{t+1}-L^{*}\rVert^{2}_{F(\Lambda_{t})}$ $\leq\left(27\sqrt{3}\gamma p\mu r\right)\frac{1}{c_{\alpha}\mu r\kappa}\frac{25}{4}\frac{\sigma_{1}^{*}}{\sigma_{r}^{*}}\|L_{t+1}-L^{*}\|_{F}^{2}$
$=\left(27\sqrt{3}p\right)\frac{25}{4\sqrt{c_{\alpha}}}\|L_{t+1}-L^{*}\|_{F}^{2}.$ (37)
Next we lower bound $\lVert L_{t+1}-L^{*}\rVert^{2}_{F(\Omega)}$ we apply Lemma C.2. Recall that by Lemma C.4, $\left(U_{t+1},V_{t+1}\right)\in\mathcal{B}_{\textnormal{err}}(\frac{13}{\sqrt{c_{e}}})\cap\mathcal{B}_{\textnormal{bln}}(\frac{1}{c_{l}})\cap\mathcal{B}_{\mu}$. Hence, for a large enough $c_{e}$, $\left(U_{t+1},V_{t+1}\right)$ satisfies the condition of Lemma C.2, with $\epsilon=\frac{1}{20}$. By Lemma C.2, w.p. at least $1-\frac{3}{n_{1}^{3}}$
$\|L_{t+1}-L^{*}\|^{2}_{F(\Omega)}\geq(1-\tfrac{1}{20})^{2}p\lVert L_{t+1}-L^{*}\rVert^{2}_{F}$ (38)
Inserting eq. 37 and eq. 38 into eq. 35 yields
$\lVert L_{t+1}-L^{*}\rVert^{2}_{F(\Omega_{t})}\geq p\left(\left(\tfrac{95}{100}\right)^{2}-\tfrac{27\sqrt{3}\cdot 25}{4\sqrt{c_{\alpha}}}\right)\|L_{t+1}-L^{*}\|_{F}^{2}\geq p\left(\tfrac{9}{10}\right)^{2}\|L_{t+1}-L^{*}\|_{F}^{2}$
where the last inequality follows for a large enough $c_{\alpha}$. We conclude that
$\|L_{t+1}-L^{*}\|_{F}\leq\frac{10}{9\sqrt{p}}\sqrt{p}\lVert L_{t+1}-L^{*}\rVert_{F(\Omega\setminus\Lambda_{t})}.$
By eq. 34 the right hand size is upper bounded by $\frac{\sigma_{r}^{*}}{2^{t+1}c_{e}\sqrt{\kappa}}$, which implies that $(U_{t+1},V_{t+1})\in\mathcal{B}_{\textnormal{err}}\left(\frac{1}{2^{t+1}c_{e}\sqrt{\kappa}}\right)$. Since $(U_{t+1},V_{t+1})\in\mathcal{B}_{\textnormal{bln}}(\frac{1}{c_{l}})\cap\mathcal{B}_{\mu}$ was proved at the beginning, $(U_{t+1},V_{t+1})$ satisfy eq. 27, which concludes the proof. ∎
C.3 Auxiliary Lemmas for Theorem 3.2
The first result is Lemma SM2.7 in (Zilber and Nadler,, 2022).
Lemma C.8.
Let $Z=\begin{pmatrix}U\\ V\end{pmatrix},Z^{\prime}=\begin{pmatrix}U^{\prime}\\ V^{\prime}\end{pmatrix}\in\mathbb{R}^{(n_{1}+n_{2})\times r}$. Denote
$a=\left(\sqrt{2}\max\{\sigma_{1}(U),\sigma_{1}(V)\}d_{P}(Z^{\prime},Z)+\tfrac{1}{2}d_{P}(Z^{\prime},Z)\right)d_{P}(Z^{\prime},Z).$ (39)
Then
$\|U^{\prime\top}U^{\prime}-V^{\prime\top}V^{\prime}\|_{F}$ $\leq\|U^{\top}U-V^{\top}V\|_{F}+2a,$ (40)
$\|U^{\prime}V^{\prime\top}-UV^{\top}\|_{F}$ $\leq a.$ (41)
The second lemma is proven by Yi et al., (2016) as part of their proof of Theorem 3. It bounds the spectral distance of an initial estimate of $L^{*}$ obtained by removing the largest-magnitude entries and projecting to rank-$r$ matrices.
Lemma C.9.
Let $X=L^{*}+S^{*}$ where $L^{*}\in\mathcal{M}(n_{1},n_{2},r,\mu)$ with singular values $\sigma_{1}^{*}\geq\dots\geq\sigma_{r}^{*}$, and $S^{*}$ satisfies Assumption 3 for some known $10<\alpha<$ 1. Let $L=UV^{\top}$ be an estimate of $L^{*}$ constructed according to the following equation,
$(U,V)=\textnormal{b-SVD}_{r}\left[\tfrac{1}{p}\left(\mathcal{P}_{\Omega}(X)-\mathcal{T}_{\alpha}\left(\mathcal{P}_{\Omega}(X),\Omega\right)\right)\right].$ (42)
There exist a constant $c$ such that if $p\geq 4\frac{\mu r^{2}\log n_{1}}{\epsilon^{2}n_{2}}$, then w.p. at least $1-\frac{6}{n_{1}}$ the spectral distance between $L$ and $L^{*}$ is bounded as follows,
$\|L-L^{*}\|_{op}\leq 16\alpha\mu r\sigma_{1}^{*}+\frac{2c\epsilon\sigma_{1}^{*}}{\sqrt{r}}.$ (43)
The third lemma is a direct consequence of Lemma SM2.4 in (Zilber and Nadler,, 2022).
Lemma C.10.
Let $L^{*}$ be a matrix of rank $r$ and let $Z^{*}=\textnormal{b-SVD}_{r}[L^{*}]$. For any $Z=\begin{pmatrix}U\\ V\end{pmatrix}$ that is perfectly balanced, namely $\|U^{\top}U-V^{\top}V\|_{F}=0$, it holds that
$d_{P}(Z,Z^{*})\leq\frac{\sqrt{2r}}{\sqrt{(\sqrt{2}-1)\sigma^{*}_{r}}}\|UV^{\top}-L^{*}\|_{op}.$ (44)
Recall the definition of $S_{init}$, eq. 16. The following lemma bounds the operator norm of $Z=\text{b-SVD}_{r}{\big{[}}\frac{1}{p}\left(\mathcal{P}_{\Omega}(X)-S_{init}\right){\big{]}}$. Its proof appears in Appendix D.4.
Lemma C.11.
Let $L^{*}\in\mathcal{M}(n_{1},n_{2},r,\mu)$ with largest singular value $\sigma_{1}^{*}$. Let $Z^{*}=\begin{pmatrix}U^{*}\\ V^{*}\end{pmatrix}=\textnormal{b-SVD}_{r}[L^{*}]$ and $Z=\begin{pmatrix}U\\ V\end{pmatrix}$ where $U,V$ are perfectly balanced. If $\|UV^{\top}-L^{*}\|_{op}\leq\frac{\sigma_{1}^{*}}{4}$ then
$\|Z\|_{op}=\sqrt{2\sigma_{1}(UV^{\top})}\leq\sqrt{\frac{5}{2}\sigma_{1}^{*}}$ (45)
In the last step of the algorithm we apply the clipping operator $R_{\eta}$, see eq. 18. The following lemma , whose proof appears in Appendix D.5, state that under some conditions applying the clipping operator with the right $\eta$ reduces the Procrustes distance.
Lemma C.12.
Let $L^{*}\in\mathcal{M}(n_{1},n_{2},r,\mu)$, $Z^{*}=\begin{pmatrix}U^{*}\\ V^{*}\end{pmatrix}=\textnormal{b-SVD}_{r}[L^{*}]$ and $Z=\begin{pmatrix}U\\ V\end{pmatrix}=\textnormal{b-SVD}_{r}[A]$ for some matrix $A\in\mathbb{R}^{n_{1}\times n_{2}}$. If $\|UV^{\top}-L^{*}\|_{op}\leq\frac{\sigma_{1}^{*}}{4}$ then for $\eta_{1}=\sqrt{\frac{\mu r}{n_{1}}}\|Z\|_{op}$, $\eta_{2}=\sqrt{\frac{\mu r}{n_{2}}}\|Z\|_{op}$ and $Z^{\prime}=\begin{pmatrix}R_{\eta_{1}}(U)\\ R_{\eta_{2}}(V)\end{pmatrix}$ it follows that
$d_{P}(Z^{\prime},Z^{*})\leq d_{P}(Z,Z^{*}).$ (46)
C.4 Proof of Theorem 3.2
Proof.
First we would prove that $L_{0}=U_{0}V_{0}^{\top}$ is a sufficiently accurate estimate of $L^{*}$ and that it is balanced. Formally we would like to bound the terms $\|U_{0}^{\top}U_{0}-V_{0}^{\top}V_{0}\|_{F}$ and $\|L_{0}-L^{*}\|_{F}$. We denote $Z^{*}=\begin{pmatrix}U^{*}\\ V^{*}\end{pmatrix}=\textnormal{b-SVD}_{r}[L^{*}]$ and recall that $Z_{0}=\begin{pmatrix}U_{0}\\ V_{0}\end{pmatrix}$. We denote
$a={\big{(}}\sqrt{2\sigma^{*}_{1}}+\frac{1}{2}d_{P}(Z_{0},Z^{*}){\big{)}}d_{P}(Z_{0},Z^{*}).$
Since $\|U^{*\top}U^{*}-V^{*\top}V^{*}\|=0$ and $U^{*}V^{*\top}=L^{*}$ by eq. 40 and eq. 41 of Lemma C.8
$\begin{split}&\|U_{0}^{\top}U_{0}-V_{0}^{\top}V_{0}\|_{F}\leq 2a,\\ &\|L_{0}-L^{*}\|_{F}\leq a.\end{split}$ (47)
To bound $a$ we need to bound $d_{P}(Z_{0},Z^{*})$. We do so by showing that by Lemma C.12 $d_{P}(Z_{0},Z^{*})\leq d_{P}(Z_{,}Z^{*})$. To apply the lemma we first prove its conditions hold. Specifically we prove that $\|UV^{\top}-L^{*}\|_{op}\leq\frac{\sigma_{1}^{*}}{4}$. Recall that $S_{init}=\mathcal{T}_{\alpha}{\left(\mathcal{P}_{\Omega}(X),\Omega\right)}$ and that $Z=\begin{pmatrix}U\\ V\end{pmatrix}=\text{b-SVD}_{r}{\big{[}}\frac{1}{p}\left(\mathcal{P}_{\Omega}(X)-S_{init}\right){\big{]}}$. Since $\Omega$ follows Assumption 2 with $p\geq C\frac{\mu r^{2}\kappa^{4}\log n_{1}}{n_{2}}$, for a large enough $C$, we can apply Lemma C.9 with $\epsilon=\frac{1}{16c}$. By Lemma C.9 w.p. at least $1-\frac{6}{n_{1}}$
$\|UV^{\top}-L^{*}\|_{op}\leq 16\alpha\mu r\sigma_{1}^{*}+\frac{2c\epsilon\sigma_{1}^{*}}{\sqrt{r}}\leq 16\alpha\mu r\sigma_{1}^{*}+\frac{\sigma_{1}^{*}}{8\sqrt{r}}.$
Since we assumed $\alpha\leq\frac{1}{c_{\alpha}\kappa^{2}r^{\frac{3}{2}}\mu}$, for a large enough $c_{\alpha}$
$\|UV^{\top}-L^{*}\|_{op}\leq\frac{16\sigma_{1}^{*}}{c_{\alpha}\kappa^{2}\sqrt{r}}+\frac{\sigma_{1}^{*}}{8\sqrt{r}}\leq\frac{\sigma_{1}^{*}}{4}.$
Since $\eta_{1}=\sqrt{\frac{\mu r}{n_{1}}}\|Z\|_{op}$, $\eta_{2}=\sqrt{\frac{\mu r}{n_{2}}}\|Z\|_{op}$ and $Z_{0}=\begin{pmatrix}R_{\eta_{1}}(U)\\ R_{\eta_{2}}(V)\end{pmatrix}$ by Lemma C.12
$d_{P}(Z_{0},Z^{*})\leq d_{P}(Z,Z^{*}).$
Hence,
$a\leq{\big{(}}\sqrt{2\sigma^{*}_{1}}+\frac{1}{2}d_{P}(Z,Z^{*}){\big{)}}d_{P}(Z,Z^{*}).$ (48)
[…]
$d_{P}(Z,Z^{*})\leq\frac{\sqrt{2r}}{\sqrt{(\sqrt{2}-1)\sigma^{*}_{r}}}\|UV^{\top}-L^{*}\|_{op}.$ (49)
To bound $\|UV^{\top}-L^{*}\|_{op}$ we apply Lemma C.9. Since $\Omega$ follows Assumption 2 with $p\geq C\frac{\mu r^{2}\kappa^{4}\log n_{1}}{n_{2}}$ we can apply Lemma C.9 with $\epsilon=\frac{2}{\sqrt{C}k^{2}}$. This gives that w.p. at least $1-\frac{6}{n_{1}^{2}}$
$d_{P}(Z,Z^{*})$ $\leq\frac{\sqrt{2r}}{\sqrt{(\sqrt{2}-1)\sigma^{*}_{r}}}\left(16\alpha\mu r\sigma_{1}^{*}+\frac{2c\epsilon\sigma_{1}^{*}}{\sqrt{r}}\right)=3\sqrt{\sigma_{r}^{*}}\left(16\alpha\mu\kappa r^{\frac{3}{2}}+{\frac{4c}{\sqrt{C}\kappa}}\right).$
Since we assumed $\alpha\leq\frac{1}{c_{\alpha}\kappa^{2}r^{\frac{3}{2}}\mu}$
$d_{P}(Z,Z^{*})\leq 3\sqrt{\sigma_{r}^{*}}\left(\frac{16}{c_{\alpha}\kappa}+\frac{4c}{\sqrt{C}\kappa}\right)\leq 3\frac{\sqrt{\sigma_{r}^{*}}}{\kappa}\left(\frac{16}{c_{\alpha}}+\frac{4c}{\sqrt{C}}\right).$ (50)
Denote $\xi=\left(\frac{16}{c_{\alpha}}+\frac{4c}{\sqrt{C}}\right)$. Inserting eq. 50 into eq. 48 gives
$a$ $\leq\left(\sqrt{2\sigma^{*}_{1}}+3\frac{\sqrt{\sigma_{r}^{*}}}{\kappa}\xi\right)3\frac{\sqrt{\sigma_{r}^{*}}}{\kappa}\xi=\frac{\sigma_{r}^{*}}{\sqrt{\kappa}}\left(3\sqrt{2}+\frac{9}{\kappa^{1.5}}\xi\right)\xi.$
Note that for any $\varepsilon>0$ for large enough $C,c_{\alpha}$ it follows that $\xi<\varepsilon$. Therefore for large enough constants $c_{\alpha}$ and $C$
$a\leq\min\left\{\frac{\sigma_{r}^{*}}{c_{e}\sqrt{\kappa}},\frac{\sigma_{r}^{*}}{4c_{l}}\right\}.$
Inserting the above bound into eq. 47 gives
$(U_{0},V_{0})\in\mathcal{B}_{\textnormal{err}}(\frac{1}{c_{e}\sqrt{\kappa}})\cap\mathcal{B}_{\textnormal{bln}}(\frac{1}{2c_{l}}).$
Next we prove that $(U_{0},V_{0})\in\mathcal{B}_{\mu}$, which requires to bound $\|U_{0}\|_{2,\infty}$ and $\|V_{0}\|_{2,\infty}$. Since $\begin{pmatrix}U_{0}\\ V_{0}\end{pmatrix}=\begin{pmatrix}R_{\eta_{1}}(U)\\ R_{\eta_{2}}(V)\end{pmatrix}$ by the definition, eq. 18, of the clipping operator $R_{\eta}$
$\|U_{0}\|_{2,\infty}\leq\eta_{1}=\sqrt{\frac{\mu r}{n}}\|Z\|_{op},\quad\|V_{0}\|_{2,\infty}\leq\eta_{2}=\sqrt{\frac{\mu r}{n_{2}}}\|Z\|_{op}.$
Since we proved that $\|UV^{\top}-L^{*}\|\leq\frac{\sigma_{1}^{*}}{4}$ we can apply Lemma C.11. Specifically inserting eq. 45 into the above inequality completes the proof as it gives
$\|U_{0}\|_{2,\infty}\leq\sqrt{\frac{\mu r}{n_{1}}}\|Z\|_{op}=\sqrt{\frac{2\mu r\sigma_{1}(UV^{\top})}{n_{1}}}\leq\sqrt{\frac{5\sigma_{1}^{*}}{2}\frac{\mu r}{n_{1}}}\leq\sqrt{\frac{3\mu r\sigma_{1}^{*}}{n_{1}}},$
$\|V_{0}\|_{2,\infty}\leq\sqrt{\frac{\mu r}{n_{2}}}\|Z\|_{op}=\sqrt{\frac{2\mu r\sigma_{1}(UV^{\top})}{n_{2}}}\leq\sqrt{\frac{5\sigma_{1}^{*}}{2}\frac{\mu r}{n_{2}}}\leq\sqrt{\frac{3\mu r\sigma_{1}^{*}}{n_{2}}}.$
Appendix D Proofs of Lemmas C.4, C.5, C.6, C.11 and C.12.
D.1 Proof of Lemma C.4
To prove the lemma we use the following auxiliary result, whose proof appears in Appendix E.1.
Lemma D.1.
Let $Z_{t}$ be the estimated factorization after $t$ steps as defined in Algorithm 2, and let $Z_{t+1}$ be the factorization at iteration $t+1$. Denote
$a_{t}=\sqrt{2}\max\{\sigma_{1}(U_{t}),\sigma_{1}(V_{t})\}d_{P}(Z_{t+1},Z_{t})+\tfrac{1}{2}d^{2}_{P}(Z_{t+1},Z_{t}).$
For large enough constants $c_{e},c_{l}$ the following holds. If $(U_{0},V_{0})\in\mathcal{B}_{\textnormal{err}}({1}/{c_{e}\sqrt{\kappa}})\cap\mathcal{B}_{\textnormal{bln}}({1}/{2c_{l}})\cap\mathcal{B}_{\mu}$ and $(U_{s},V_{s})\in\mathcal{C}(U_{s-1},V_{s-1},\frac{\sigma_{r}^{*}}{4^{s}c_{e}\kappa})$ for every $11\leq s\leq t+$ 1 then
$a_{t}\leq\frac{4\sqrt{2}\sigma_{r}^{*}}{2^{t+1}\sqrt{c_{e}}}+{\frac{\sigma_{r}^{*}}{2\cdot 4^{t+1}c_{e}\kappa}}.$ (51)
Proof of Lemma C.4.
We start by proving that $(U_{t+1},V_{t+1})\in\mathcal{B}_{\textnormal{bln}}(\frac{1}{c_{l}})$. By eq. 40 of Lemma C.8,
$\|U_{t+1}^{\top}U_{t+1}-V_{t+1}^{\top}V_{t+1}\|\leq\|U_{t}^{\top}U_{t}-V_{t}^{\top}V_{t}\|+2a_{t}.$
Therefore, by induction
$\|U_{t+1}^{\top}U_{t+1}-V_{t+1}^{\top}V_{t+1}\|\leq\|U_{0}^{\top}U_{0}-V_{0}^{\top}V_{0}\|+2\sum_{k=0}^{t}a_{k}.$ (52)
By eq. 51 of Lemma D.1, for large enough $c_{e}$, for each iteration step $0\leq k\leq t$
$a_{k}\leq\frac{\sigma_{r}^{*}}{2^{k+3}c_{l}}.$ (53)
Inserting eq. 53 into eq. 52 gives,
$\begin{split}\|U_{t+1}^{\top}U_{t+1}-V_{t+1}^{\top}V_{t+1}\|&\leq\|U_{0}^{\top}U_{0}-V_{0}^{\top}V_{0}\|+\frac{\sigma_{r}^{*}}{c_{l}}\sum_{k=0}^{t}\frac{1}{2^{k+2}}\\ &\leq\|U_{0}^{\top}U_{0}-V_{0}^{\top}V_{0}\|+\frac{\sigma_{r}^{*}}{2c_{l}}.\end{split}$
Since $(U_{0},V_{0})\in\mathcal{B}_{\textnormal{bln}}({1}/{2c_{l}})$ the above inequality implies that $(U_{t+1},V_{t+1})\in\mathcal{B}_{\textnormal{bln}}({1}/{c_{l}})$.
Next, we prove $(U_{t+1},V_{t+1})\in\mathcal{B}_{\textnormal{err}}(\frac{13}{\sqrt{c_{e}}})$. By the triangle inequality
$\|U_{t+1}V_{t+1}^{\top}-L^{*}\|_{F}\leq\|U_{t+1}V_{t+1}^{\top}-U_{t}V_{t}^{\top}\|_{F}+\|U_{t}V_{t}^{\top}-L^{*}\|_{F}.$
By Lemma C.8 $\|U_{t+1}V_{t+1}^{\top}-U_{t}V_{t}^{\top}\|_{F}\leq a_{t}$ and therefore,
$\|U_{t+1}V_{t+1}^{\top}-L^{*}\|_{F}\leq a_{t}+\|U_{t}V_{t}^{\top}-L^{*}\|_{F}.$
By induction it follows that
$\|U_{t+1}V_{t+1}^{\top}-L^{*}\|_{F}\leq\|U_{0}V_{0}^{\top}-L^{*}\|_{F}+\sum_{k=0}^{t}a_{k}.$ (54)
Next we use D.1 to bound $a_{k}$. Note that for $c_{e}\geq 1$ since $\kappa\geq 1$,
$a_{k}\leq\frac{4\sqrt{2}\sigma_{r}^{*}}{2^{k+1}\sqrt{c_{e}}}+\frac{\sigma_{r}^{*}}{2\cdot 4^{k+1}\sqrt{c_{e}}}\leq\frac{4\sqrt{2}\sigma_{r}^{*}}{2^{k+1}\sqrt{c_{e}}}+\frac{4\sqrt{2}\sigma_{r}^{*}}{2^{k+1}\sqrt{c_{e}}}\leq\frac{4\sqrt{2}\sigma_{r}^{*}}{2^{k}\sqrt{c_{e}}}.$
As a result of the above inequality
$\sum_{k=0}^{t}a_{k}\leq\frac{4\sqrt{2}\sigma_{r}^{*}}{\sqrt{c_{e}}}\sum_{k=0}^{t}\frac{1}{2^{k}}\leq\frac{8\sqrt{2}\sigma_{r}^{*}}{\sqrt{c_{e}}}.$ (55)
By the assumption $(U_{0},V_{0})\in\mathcal{B}_{\textnormal{err}}\left(\frac{1}{c_{e}\sqrt{\kappa}}\right)$ it follows that $\|U_{0}V_{0}^{\top}-L^{*}\|\leq\frac{\sigma_{r}^{*}}{c_{e}\sqrt{\kappa}}$. Inserting this inequality and eq. 55 into eq. 54 gives
$\|U_{t+1}V_{t+1}^{\top}-L^{*}\|_{F}\leq\frac{\sigma_{r}^{*}}{c_{e}\sqrt{\kappa}}+\frac{(8\sqrt{2})\sigma_{r}^{*}}{\sqrt{c_{e}}}\leq\frac{13\sigma^{*}_{r}}{\sqrt{c_{e}}}.$
We conclude that $(U_{t+1},V_{t+1})\in\mathcal{B}_{\textnormal{err}}(\frac{13}{\sqrt{c_{e}}})\cap\mathcal{B}_{\textnormal{bln}}(\frac{1}{c_{l}})$. ∎
D.2 Proof of Lemma C.5
To prove the Lemma C.5 we use the following four auxiliary lemmas, whose proofs are in Appendix E.2. All of them consider the following setting. Let $L_{t}=U_{t}V_{t}^{\top}$ and $\Lambda_{t}$ be the estimates of $L^{*}$ and $\Lambda_{*}$ at iteration $t$, respectively. Let $L_{t+1}=U_{t+1}V_{t+1}^{\top}$ be the updated estimate, where $(U_{t+1},V_{t+1})$ are computed by (10) with some $\delta>0$. Denote $\Delta U_{t}^{*}=U_{t}-U^{*}$, $\Delta U_{t+1}=U_{t+1}-U_{t}$ with similar definitions for $\Delta V_{t}^{*}$ and $\Delta V_{t+1}$. Denote by $\Omega_{t}=\Omega\setminus\Lambda_{t}$ the set of non removed entries, by $I_{t}^{\text{in}}=\Omega_{t}\cap\Lambda_{*}^{c}$ the set of non-removed entries which are true inliers and by $I_{t}^{\text{out}}=\Omega_{t}\cap\Lambda_{*}$ the set of non-removed entries which are outliers. Finally, recall that for factor matrices $(U,V)$ and a subset of entries $\Lambda$ the objective function value is
$\mathcal{L}^{t}_{{\Lambda}}(U,V)=\arg\min_{U,V}\|U_{t}V^{\top}+UV_{t}^{\top}-U_{t}V_{t}^{\top}-X\|_{F(\Lambda)}.$
Lemma D.2.
The error restricted to the set $I_{t}^{\text{in}}$ is bounded as follows,
$\|L_{t+1}-L^{*}\|_{F(I_{t}^{\text{in}})}\leq\mathcal{L}^{t}_{I_{t}^{\text{in}}}(U_{t+1},V_{t+1})+\|\Delta U_{t+1}\Delta V_{t+1}^{\top}\|_{F(I_{t}^{\text{in}})}.$ (56)
To upper bound the first term on the RHS of eq. 56 we use the following lemma.
Lemma D.3.
Assume that $L_{t}$ is sufficiently close to $L^{*}$ in the sense that the set $\mathcal{A}_{t}=\mathcal{B}^{*}\cap\mathcal{B}_{\mu}\cap\mathcal{C}\left(U_{t},V_{t},\frac{\delta}{4^{t+1}}\right)$ is non empty. Then
$\mathcal{L}^{t}_{I_{t}^{\text{in}}}(U_{t+1},V_{t+1})\leq\sqrt{2}{\mathcal{L}^{t}_{\Omega_{t}}(U^{*},V^{*})}-{\mathcal{L}^{t}_{I_{t}^{\text{out}}}(U_{t+1},V_{t+1})},$ (57)
for any $(U^{*},V^{*})\in\mathcal{A}_{t}$, an exact factorization of $L^{*}$.
To upper bound the RHS of eq. 57 we employ the following two lemmas. The first lemma upper bounds the first term and the second lemma lower bounds the second term.
Lemma D.4.
For any $(U^{*},V^{*})\in\mathcal{B}^{*}$
$\mathcal{L}^{t}_{\Omega_{t}}(U^{*},V^{*})\leq\|\Delta U^{*}_{t}\Delta V_{t}^{*\top}\|_{F({\Omega_{t}})}+\|S^{*}\|_{F({\Omega_{t}})}.$ (58)
Lemma D.5.
Under the above assumptions,
$\mathcal{L}^{t}_{I_{t}^{\text{out}}}(U_{t+1},V_{t+1})\geq$ $\|L_{t+1}-L^{*}-S^{*}\|_{F(I_{t}^{\text{out}})}-\|\Delta U_{t+1}\Delta V_{t+1}^{\top}\|_{F(I_{t}^{\text{out}})}.$ (59)
Proof of Lemma C.5.
Our goal is to bound $\|L_{t+1}-L^{*}\|_{F(\Omega_{t})}$. Since $\Omega_{t}=I_{t}^{\text{in}}\cup I_{t}^{\text{out}}$ ,
$\|L_{t+1}-L^{*}\|_{F(\Omega_{t})}\leq\|L_{t+1}-L^{*}\|_{F(I_{t}^{\text{in}})}+\|L_{t+1}-L^{*}\|_{F({I_{t}^{\text{out}}})}.$
Lemma D.2 bounds $\|L_{t+1}-L^{*}\|_{F(I_{t}^{\text{in}})}$,
$\|L_{t+1}-L^{*}\|_{F(\Omega_{t})}\leq$ $\mathcal{L}^{t}_{I_{t}^{\text{in}}}(U_{t+1},V_{t+1})+\|\Delta U_{t+1}\Delta V_{t+1}^{\top}\|_{F(I_{t}^{\text{in}})}+\|L_{t+1}-L^{*}\|_{F({I_{t}^{\text{out}}})}.$ (60)
To upper bound $\mathcal{L}^{t}_{I_{t}^{\text{in}}}(U_{t+1},V_{t+1})$ we insert eq. 59 and eq. 58 into eq. 57 which yields,
$\begin{split}\mathcal{L}^{t}_{I_{t}^{\text{in}}}(U_{t+1},V_{t+1})\leq&\sqrt{2}\|\Delta U_{t}^{*}\Delta V_{t}^{*\top}\|_{F({\Omega_{t}})}+\sqrt{2}\|S^{*}\|_{F({\Omega_{t}})}\\ &+\|\Delta U_{t+1}\Delta V_{t+1}^{\top}\|_{F(I_{t}^{\text{out}})}-\|L_{t+1}-L^{*}-S^{*}\|_{F(I_{t}^{\text{out}})}.\end{split}$ (61)
Inserting eq. 61 into eq. 60 and using the fact that $\|S\|_{F({\Omega_{t}})}=\|S^{*}\|_{F(I_{t}^{\text{out}})}$ gives,
$\begin{split}\|L_{t+1}-L^{*}\|_{F(\Omega_{t})}\leq&\sqrt{2}\|\Delta U_{t}^{*}\Delta V_{t}^{*\top}\|_{F({\Omega_{t}})}+\sqrt{2}\|S^{*}\|_{F(I_{t}^{\text{out}})}\\ &+\underbrace{\|\Delta U_{t+1}\Delta V_{t+1}^{\top}\|_{F(I_{t}^{\text{out}})}+{\|\Delta U_{t+1}\Delta V_{t+1}^{\top}\|_{F(I_{t}^{\text{in}})}}}_{\Delta}\\ &+\|L_{t+1}-L^{*}\|_{F({I_{t}^{\text{out}}})}-\|L_{t+1}-L^{*}-S^{*}\|_{F(I_{t}^{\text{out}})}.\end{split}$ (62)
To bound the term $\Delta$ above we note that,
$\|\Delta U_{t+1}\Delta V_{t+1}^{\top}\|^{2}_{F(I_{t}^{\text{out}})}+\|\Delta U_{t+1}\Delta V_{t+1}^{\top}\|^{2}_{F(I_{t}^{\text{in}})}=\|\Delta U_{t+1}\Delta V_{t+1}^{\top}\|^{2}_{F({\Omega_{t})}}.$
Using the inequality $a+b\leq\sqrt{2}\sqrt{a^{2}+b^{2}}$ yields
$\Delta\leq\sqrt{2}\|\Delta U_{t+1}\Delta V_{t+1}^{\top}\|_{F({\Omega_{t})}}.$ (63)
By the triangle inequality
$\|L_{t+1}-L^{*}\|_{F({I_{t}^{\text{out}}})}-\|L_{t+1}-L^{*}-S^{*}\|_{F(I_{t}^{\text{out}})}\leq\|S^{*}\|_{F(I_{t}^{\text{out}})}.$ (64)
Inserting eq. 63 and eq. 64 into eq. 62 gives,
$\|L_{t+1}-L^{*}\|_{F(\Omega_{t})}\leq$ $\sqrt{2}\left[\|\Delta U_{t}^{*}\Delta V_{t}^{*\top}\|_{F({\Omega_{t}})}+\|\Delta U_{t+1}\Delta V_{t+1}^{\top}\|_{F({\Omega_{t})}}\right]$
$+(1+\sqrt{2})\|S^{*}\|_{F(I_{t}^{\text{out}})}.$
D.3 Proofs of Lemmas C.6 and C.7
Before we prove the lemmas we present three auxiliary lemmas. The first is Lemma 10 in (Yi et al.,, 2016). It establishes that if the sampling probability $p$ is sufficiently high, then the number of sampled entries in all rows and columns of $\Omega$ concentrates around their expected value.
Lemma D.6.
Let $\Omega\subset[n_{1}]\times[n_{2}]$ with $n_{1}\geq n_{2}$ satisfy Assumption 1, namely its entries are sampled independently with probability $p$. There exist a constant $C$ such that if $p\geq C\frac{\log n_{1}}{n_{2}}$ then w.p. at least $1-\frac{6}{n_{1}}$ uniformally over all $i\in[n_{1}]$ and $j,\in[n_{2}]$
$\left|r_{i}-pn_{2}\right|\leq\frac{1}{2}pn_{2},\quad\left|c_{j}-pn_{1}\right|\leq\frac{1}{2}pn_{1}.$ (65)
The second result is Lemma 14 in (Yi et al.,, 2016) . It provides an upper bound on the error $\|UV^{\top}-U^{*}V^{*\top}\|_{F(\Lambda)}$ for a sufficiently well spread out subset $\Lambda$ of the entries.
Lemma D.7.
Let $L^{*}$ have rank $r$, largest singular value $\sigma_{1}^{*}$ and decomposition $(U^{*},V^{*})\in\mathcal{B}_{*}\cap\mathcal{B}_{\mu}$. Let $U\in\mathbb{R}^{n_{1}\times r},V\in\mathbb{R}^{n_{2}\times r}$ be a pair of matrices, such that $(U,V)\in\mathcal{B}_{\mu}$, and let $\theta\in(0,1)$. Assume that $\Lambda\subseteq[n_{1}]\times[n_{2}]$ satisfies $|\Lambda_{(i,\cdot)}|\leq\theta n_{2}$ for all $i\in[n_{1}]$ and $|\Lambda_{(\cdot,j)}|\leq\theta n_{1}$ for all $j\in[n_{2}]$ then
$\|UV^{\top}-U^{*}V^{*\top}\|^{2}_{F(\Lambda)}\leq 18\sqrt{3}\theta r\mu\sigma_{1}^{*}(\|U-U^{*}\|_{F}^{2}+\|V-V^{*}\|_{F}^{2}).$ (66)
The third lemma bounds the impact of the entries in $(\Omega\setminus\Lambda_{t})\cap\Lambda_{*}$ on the error term $\|L_{t}-X\rVert_{F(\Omega)}$. The lemma is based on a work presented in (Yi et al.,, 2016) (as part of their proof of Lemma 2), its proof appears in Appendix E.3.
Lemma D.8.
Let $X=L^{*}+S^{*}$, and $\Omega\subset[n_{1}]\times[n_{2}]$, where the corruption matrix $S^{*}$ satisfies Assumption 3 for some known $10<\alpha<$ 1. Suppose that for a given estimate $L$ of $L^{*}$, the set of corrupted entries is estimated by $\Lambda=\text{support}\left(\mathcal{T}_{\gamma\alpha}\left(L-X,\Omega\right)\right)$ with an over remvoal factor $1<\gamma\leq\frac{1}{\alpha}$. Then the error on the remaining corrupted entries $(\Omega\setminus\Lambda)\cap\Lambda^{*}$, where $\Lambda^{*}=\mbox{support}(S^{*})$, is bounded as follows,
$\|L-X\rVert_{F({(\Omega\setminus\Lambda)\cap\Lambda_{*}})}\leq\sqrt{\frac{2}{\gamma-1}}\lVert L-L^{*}\rVert_{F(\Omega)}.$ (67)
Proof of Lemma C.6.
By the triangle inequality
$\|S^{*}\|_{F((\Omega\setminus\Lambda)\cap\Lambda_{*})}$ $=\|L-(L^{*}+S^{*})-(L-L^{*})\|_{F((\Omega\setminus\Lambda)\cap\Lambda_{*})}$
$\leq\underbrace{\|L-L^{*}\|_{F((\Omega\setminus\Lambda)\cap\Lambda_{*})}}_{W_{1}}+\underbrace{\|(L-L^{*})-S^{*}\|_{F((\Omega\setminus\Lambda)\cap\Lambda_{*})}}_{W_{2}}.$ (68)
First we bound $W_{1}$. To this end we note that since $\Omega$ follows Assumption 2 with $p\geq C\frac{\log n_{1}}{n_{2}}$ by Lemma D.6 w.p. at least $1-\frac{6}{n_{1}}$ for every $i\in[n_{1}]$ and $j\in[n_{2}]$
$\left|r_{i}-pn_{2}\right|\leq\frac{1}{2}pn_{2},\quad\left|c_{j}-pn_{1}\right|\leq\frac{1}{2}pn_{1}.$
By Assumption 3, $S^{*}\in\mathcal{S}_{\alpha}^{\Omega}$ and therefore
$|(\Lambda_{*})_{(i,\cdot)}|\leq\alpha r_{i},\quad|(\Lambda_{*})_{(\cdot,j)}|\leq\alpha c_{j},$
Since $(\Omega\setminus\Lambda_{t})\cap\Lambda_{*}\subset\Lambda_{*}$
$|((\Omega\setminus\Lambda_{t})\cap\Lambda_{*})_{(i,\cdot)}|\leq|(\Lambda_{*})_{(i,\cdot)}|\leq\alpha r_{i}\leq\frac{3}{2}\alpha pn_{2}$
and similarly $|(\Lambda_{*})_{(\cdot,j)}|\leq\frac{3}{2}\alpha pn_{1}$. Therefore we can invoke Lemma D.7 with $\theta=\frac{3}{2}\alpha p$ to obtain the following bound on $W_{1}$,
$W_{1}^{2}={\|UV^{\top}-U^{*}V^{*\top}\|^{2}_{F((\Omega\setminus\Lambda)\cap\Lambda_{*})}}\leq 27\sqrt{3}\alpha pr\mu\sigma_{1}^{*}(\|U-U^{*}\|_{F}^{2}+\|V-V^{*}\|_{F}^{2}).$
Since $(U^{*},V^{*})$ satisfy eq. 24 it follows that
$W_{1}\leq\sqrt{27\sqrt{3}\alpha pr\mu\sigma_{1}^{*}\frac{25}{4\sigma_{r}^{*}}\|L-L\|_{F}^{2}}\leq{18\sqrt{\alpha p\mu r\kappa}}\lVert L-L^{*}\rVert_{F}.$
Next we bound $W_{2}$ in eq. 68. To this end we invoke Lemma D.8 which gives,
$W_{2}={\|L_{t}-L^{*}-S^{*}\|_{F((\Omega\setminus\Lambda_{t})\cap\Lambda_{*})}}\leq\sqrt{\frac{2}{\gamma-1}}\lVert L_{t}-L^{*}\rVert_{F(\Omega)}.$
Inserting the bounds on $W_{1}$ and $W_{2}$ into eq. 68 yields eq. 25. ∎
Proof of Lemma C.7.
Since $\Omega$ follows Assumption 2 with $p\geq C\frac{\log n_{1}}{n_{2}}$ by Lemma D.6 w.p. at least $1-\frac{6}{n_{1}}$ for every $i\in[n_{1}]$ and $j\in[n_{2}]$
$\left|r_{i}-pn_{2}\right|\leq\frac{1}{2}pn_{2},\quad\left|c_{j}-pn_{1}\right|\leq\frac{1}{2}pn_{1}.$
By the definition of $\mathcal{T}_{\gamma\alpha}$, eq. 6, for any matrix $A$, it follows that if $\Lambda=\text{support}\left(\mathcal{T}_{\gamma\alpha}\left(A,\Omega\right)\right)$ then
$|(\Lambda)_{(i,\cdot)}|\leq\gamma\alpha r_{i}\leq\frac{3}{2}\gamma\alpha pn_{2},\quad|(\Lambda)_{(\cdot,j)}|\leq\gamma\alpha c_{j}\leq\frac{3}{2}\gamma\alpha pn_{1}.$
Using Lemma D.7 with $\Lambda$ and $\theta=\frac{3}{2}\gamma\alpha p$ we obtain eq. 26. ∎
D.4 Proof of Lemma C.11
Proof.
First we bound $\sigma_{1}(UV^{\top})$. By the assumption $\|UV^{\top}-L^{*}\|_{op}\leq\frac{\sigma_{1}^{*}}{4}$ and the triangle inequality,
$\sigma_{1}(UV^{\top})\leq\frac{5}{4}\sigma_{1}^{*}.$
Next we bound $\|Z\|_{op}$. Since $U,V$ are perfectly balanced, there exists $\tilde{U}\in\mathbb{R}^{n_{1}\times r},\tilde{V}\in\mathbb{R}^{n_{2}\times r}$ unitary matrices and $\tilde{\Sigma}\in\mathbb{R}^{r\times r}$ a diagonal matrix such that
$U=\tilde{U}\tilde{\Sigma}^{\frac{1}{2}},\quad\mbox{and}\quad V=\tilde{V}\tilde{\Sigma}^{\frac{1}{2}}.$
Note that $UV^{\top}=\tilde{U}\tilde{\Sigma}\tilde{V}$, and thus $\sigma_{1}(UV^{\top})=\lambda_{1}(\tilde{\Sigma})$. In addition, $Z^{\top}Z=U^{\top}U+V^{\top}V=2\tilde{\Sigma}$. Hence,
$\|Z\|_{op}=\sqrt{\lambda_{1}(Z^{\top}Z)}=\sqrt{\lambda_{1}(2\tilde{\Sigma})}=\sqrt{2\sigma_{1}(UV^{\top})}\leq\sqrt{\frac{5\sigma_{1}^{*}}{2}}.$
D.5 Proof of Lemma C.12
To prove the lemma we use the following auxiliary lemma which is Lemma 11 in (Zheng and Lafferty,, 2016).
Lemma D.9.
Let $y\in\mathbb{R}^{r}$ be a vector such that $\|{y}\|\leq\eta$. Then for any $x\in\mathbb{R}^{r}$
$\|{R_{\eta}}(x)-y\|^{2}\leq\|{x-y}\|^{2}.$ (69)
proof of Lemma C.12.
Recall that
$d_{p}(Z^{\prime},Z^{*})=\min_{P}\|Z^{\prime}-Z^{*}P\|_{F},\quad d_{p}(Z,Z^{*})=\min_{P}\|Z-Z^{*}P\|_{F}.$
where $P$ is orthogonal. In what follows we prove that for any orthogonal matrix $\|Z^{\prime}-Z^{*}P\|_{F}\leq\|Z-Z^{*}P\|_{F}$. Let $P$ be an orthogonal matrix,
$\|{Z^{\prime}-Z^{*}P}\|_{F}^{2}=\sum_{i=1}^{n_{1}}\|R_{\eta_{1}}(U_{(i,\cdot)})-(U^{*}P)_{(i,\cdot)}\|^{2}+\sum_{i=n_{1}+1}^{n_{1}+n_{2}}\|R_{\eta_{2}}(V_{(i,\cdot)})-(V^{*}P)_{(i,\cdot)}\|^{2},$ (70)
In order to bound $\|R_{\eta_{1}}(U_{(i,\cdot)})-(U^{*}P)_{(i,\cdot)}\|^{2}$ and $\|R_{\eta_{2}}(V_{(i,\cdot)})-(V^{*}P)_{(i,\cdot)}\|^{2}$ using Lemma D.9 we first need to prove that $\|(U^{*}P)_{(i,\cdot)}\|\leq\eta_{1},\|(V^{*}P)_{(i,\cdot)}\|\leq\eta_{2}$. Since $L^{*}$ has an incoherence parameter of $\mu$
$\|U^{*}\|_{2,\infty}\leq\sqrt{\frac{r\mu\sigma_{1}^{*}}{n_{1}}},\quad\|V^{*}\|_{2,\infty}\leq\sqrt{\frac{r\mu\sigma_{1}^{*}}{n_{2}}}.$
Since we assumed $\|UV^{\top}-L^{*}\|_{op}\leq\frac{\sigma_{1}^{*}}{4}$, by the triangle inequality $\sigma_{1}^{*}\leq 2\sigma_{1}(UV^{\top})$. Therefore
$\|U^{*}\|_{2,\infty}\leq\sqrt{\frac{r\mu\sigma_{1}^{*}}{n_{1}}}\leq\sqrt{\frac{2r\mu\sigma_{1}(UV^{\top})}{n_{1}}},$
$\|V^{*}\|_{2,\infty}\leq\sqrt{\frac{r\mu\sigma_{1}^{*}}{n_{2}}}\leq\sqrt{\frac{2r\mu\sigma_{1}(UV^{\top})}{n_{2}}}.$
By Lemma C.11 $\sqrt{2\sigma_{1}(UV^{\top})}=\|Z\|_{op}$, hence
$\|U^{*}\|_{2,\infty}\leq\sqrt{\frac{r\mu}{n_{1}}}\|Z\|_{op}=\eta_{1},$
$\|V^{*}\|_{2,\infty}\leq\sqrt{\frac{r\mu}{n_{2}}}\|Z\|_{op}=\eta_{2}.$
Let $P$ be an orthogonal matrix then
$\|U^{*}P\|_{2,\infty}\leq\|U^{*}\|_{2,\infty}\|P^{\top}\|_{op}=\|U^{*}\|_{2,\infty},$
$\|V^{*}P\|_{2,\infty}\leq\|V^{*}\|_{2,\infty}\|P^{\top}\|_{op}=\|V^{*}\|_{2,\infty}.$
Therefore, $\|(U^{*}P)_{(i,\cdot)}\|\leq\eta_{1},\|(V^{*}P)_{(i,\cdot)}\|\leq\eta_{2}$. By Lemma D.9
$\|R_{\eta_{1}}(U_{(i,\cdot)})-(U^{*}P)_{(i,\cdot)}\|^{2}\leq\|{U_{(i,\cdot)}-(U^{*}P)_{(i,\cdot)}}\|^{2},$
and a similar result follows for $V^{*}$ with $\eta_{2}$. Inserting these bounds into eq. 70 gives
$\|Z^{\prime}-Z^{*}P\|_{F}^{2}\leq\sum_{i=1}^{n_{1}}\|U_{(i,\cdot)}-(U^{*}P)_{(i,\cdot)}\|^{2}+\sum_{i=n_{1}+1}^{n_{1}+n_{2}}\|V_{(i,\cdot)}-(V^{*}P)_{(i,\cdot)}\|^{2}=\|Z-Z^{*}P\|_{F}^{2}.$
Since this is true for any orthogonal matrix it follows that
$d_{P}(Z^{\prime},Z^{*})$ $\leq d_{P}(Z,Z^{*}).$
Appendix E Proofs of Lemmas D.1, D.2, D.3 D.5, D.4 and D.8.
E.1 Proof of Lemma D.1.
As part of the proof we use the following lemma, whose proof appears in Appendix F.
Lemma E.1.
There exist constants $c_{e},c_{l}$ such that the following holds. If $(U_{0},V_{0})\in\mathcal{B}_{\textnormal{err}}\left(\frac{1}{c_{e}\sqrt{\kappa}}\right)\cap\mathcal{B}_{\textnormal{bln}}\left(\frac{1}{2c_{l}}\right)\cap\mathcal{B}_{\mu}$ then
$\max\{\sigma_{1}(U_{0}),\sigma_{1}(V_{0})\}\leq 2\sqrt{\sigma_{1}^{*}}.$ (71)
Additionally if for every $11\leq s\leq t+$ 1, $(U_{s},V_{s})\in\mathcal{C}(U_{s-1},V_{s-1},\frac{\sigma_{r}^{*}}{4^{s}c_{e}\sqrt{\kappa}})$ then
$\max\{\sigma_{1}(U_{t+1}),\sigma_{1}(V_{t+1})\}\leq 4\sqrt{\sigma_{1}^{*}}.$ (72)
Proof of Lemma D.1.
Recall that
$a_{t}=\sqrt{2}\max\{\sigma_{1}(U_{t}),\sigma_{1}(V_{t})\}d_{P}(Z_{t+1},Z_{t})+\tfrac{1}{2}d^{2}_{P}(Z_{t+1},Z_{t}).$
Since for every $1\leq s\leq t$, $(U_{s},V_{s})\in\mathcal{C}(U_{s-1},V_{s-1},\frac{\sigma_{r}^{*}}{4^{s}c_{e}\sqrt{\kappa}})$ by Lemma E.1 it follows that $\max\{\sigma_{1}(U_{t}),\sigma_{1}(V_{t})\}\leq 4\sqrt{\sigma_{1}^{*}}$. Therefore
$a_{t}\leq\left(\sqrt{32}\sqrt{\sigma_{1}^{*}}+\tfrac{1}{2}d_{P}(Z_{t},Z_{t+1})\right)d_{P}(Z_{t},Z_{t+1}).$
Since $d_{P}(Z_{t},Z_{t+1})\leq\|Z_{t}-Z_{t+1}\|_{F}=\sqrt{\|U_{t+1}-U_{t}\|_{F}^{2}+\|V_{t+1}-V_{t}\|_{F}^{2}}$ we can bound $a_{t}$ as follows,
$a_{t}\leq\left(\sqrt{32\left(\|U_{t+1}-U_{t}\|_{F}^{2}+\|V_{t+1}-V_{t}\|_{F}^{2}\right)}\sqrt{\sigma_{1}^{*}}+\tfrac{1}{2}\|U_{t+1}-U_{t}\|_{F}^{2}+\|V_{t+1}-V_{t}\|_{F}^{2}\right).$
Since $\left(U_{t+1},V_{t+1}\right)\in\mathcal{C}\left(U_{t},V_{t},\frac{\sigma_{r}^{*}}{4^{t+1}c_{e}\kappa}\right)$, see eq. 8, $\|U_{t+1}-U_{t}\|_{F}^{2}+\|V_{t+1}-V_{t}\|_{F}^{2}\leq\frac{\sigma_{r}^{*}}{4^{t+1}c_{e}\kappa}$. Hence,
$\begin{split}a_{t}&\leq\frac{\sqrt{32\sigma_{r}^{*}\sigma_{1}^{*}}}{2^{t+1}\sqrt{\kappa c_{e}}}+{\frac{\sigma_{r}^{*}}{2\cdot 4^{t+1}c_{e}\kappa}}=\frac{4\sqrt{2}\sigma_{r}^{*}}{2^{t+1}\sqrt{c_{e}}}+{\frac{\sigma_{r}^{*}}{2\cdot 4^{t+1}c_{e}\kappa}}.\end{split}$
E.2 Proofs of Lemmas D.2, D.3 D.5 and D.4
Proof of Lemma D.2.
We first lower bound $\mathcal{L}^{t}_{I_{t}^{\text{in}}}(U_{t+1},V_{t+1})$, the analogue of the objective function in eq. 1 on the set $I_{t}^{\text{in}}$. By definition $S^{*}_{i,j}=0$ for every $(i,j)\in\Lambda_{*}^{c}$. Since $I_{t}^{\text{in}}\subseteq\Lambda_{*}^{c}$
$\mathcal{L}^{t}_{I_{t}^{\text{in}}}(U_{t+1},V_{t+1})$ $=\|U_{t}V_{t+1}^{\top}+U_{t+1}V_{t}^{\top}-U_{t}V_{t}^{\top}-L^{*}-S^{*}\|_{F(I_{t}^{\text{in}})}$
$=\|U_{t}V_{t+1}^{\top}+U_{t+1}V_{t}^{\top}-U_{t}V_{t}^{\top}-L^{*}\|_{F(I_{t}^{\text{in}})}.$ (73)
Note that
$U_{t}V_{t+1}^{\top}+U_{t+1}V_{t}^{\top}-U_{t}V_{t}^{\top}-L^{*}=$ $\left(U_{t}V_{t+1}^{\top}+U_{t+1}V_{t}^{\top}-U_{t}V_{t}^{\top}-U_{t+1}V_{t+1}^{\top}\right)$
$+\left(U_{t+1}V_{t+1}^{\top}-L^{*}\right)$
$=$ $(L_{t+1}-L^{*})-(U_{t+1}-U_{t})(V_{t+1}-V_{t})^{\top}.$
Inserting this into eq. 73, by the triangle inequality
$\begin{split}\mathcal{L}^{t}_{I_{t}^{\text{in}}}(U_{t+1},V_{t+1})\geq\|L_{t+1}-L^{*}\|_{F(I_{t}^{\text{in}})}-\|\Delta U_{t+1}\Delta V_{t+1}^{\top}\|_{F(I_{t}^{\text{in}})}.\end{split}$
Rearranging the terms yields,
$\|L_{t+1}-L^{*}\|_{F(\Omega_{t})}\leq\mathcal{L}^{t}_{I_{t}^{\text{in}}}(U_{t+1},V_{t+1})+\|\Delta U_{t+1}\Delta V_{t+1}^{\top}\|_{F(I_{t}^{\text{in}})}.$
Proof of Lemma D.3.
Our goal is to upper bound $\mathcal{L}^{t}_{I_{t}^{\text{in}}}(U_{t+1},V_{t+1})$. Since $\Omega\setminus\Lambda_{t}=I_{t}^{\text{in}}\cup I_{t}^{\text{out}}$
$\left[\mathcal{L}^{t}_{\Omega_{t}}(U_{t+1},V_{t+1})\right]^{2}=\left[\mathcal{L}^{t}_{I_{t}^{\text{in}}}(U_{t+1},V_{t+1})\right]^{2}+\left[\mathcal{L}^{t}_{I_{t}^{\text{out}}}(U_{t+1},V_{t+1})\right]^{2}.$
Since $\sqrt{a^{2}+b^{2}}\geq\frac{1}{\sqrt{2}}\left(a+b\right)$
$\mathcal{L}^{t}_{\Omega_{t}}(U_{t+1},V_{t+1})\geq\frac{1}{\sqrt{2}}[\mathcal{L}^{t}_{I_{t}^{\text{in}}}(U_{t+1},V_{t+1})+\mathcal{L}^{t}_{I_{t}^{\text{out}}}(U_{t+1},V_{t+1})].$
Rearranging the terms yields,
$\mathcal{L}^{t}_{I_{t}^{\text{in}}}(U_{t+1},V_{t+1})\leq\sqrt{2}\mathcal{L}^{t}_{\Omega_{t}}(U_{t+1},V_{t+1})-\mathcal{L}^{t}_{I_{t}^{\text{out}}}(U_{t+1},V_{t+1}).$ (74)
Note that $(U_{t+1},V_{t+1})=\arg\min\{\mathcal{L}^{t}_{\Omega_{t}}(U,V)\mid{(U,V)\in\mathcal{B}_{\mu}\cap\mathcal{C}\left(U_{t},V_{t},\frac{\delta}{4^{t+1}}\right)}\}$. By definition $\mathcal{A}_{t}\subset\mathcal{B}_{\mu}\cap\mathcal{C}\left(U_{t},V_{t},\frac{\delta}{4^{t+1}}\right)$. Therefore for any $(U^{*},V^{*})\in\mathcal{A}_{t}$,
$\mathcal{L}^{t}_{\Omega_{t}}(U_{t+1},V_{t+1})\leq\mathcal{L}^{t}_{\Omega_{t}}(U^{*},V^{*}).$ (75)
Inserting eq. 75 into eq. 74 yields eq. 57 of the lemma. ∎
Proof of Lemma D.4.
By the definition of $\mathcal{L}^{t}_{\Omega_{t}}$, see eq. 1, since $X=L^{*}+S^{*}$
$\mathcal{L}^{t}_{\Omega_{t}}(U^{*},V^{*})=\|U_{t}V^{*\top}+U^{*}V_{t}^{\top}-U_{t}V_{t}^{\top}-{L^{*}}-S^{*}\|_{F(\Omega_{t})}.$
Combining the triangle inequality with the fact that $L^{*}=U^{*}V^{*\top}$ gives
$\mathcal{L}^{t}_{\Omega_{t}}(U^{*},V^{*})$ $\leq\|U_{t}V^{*\top}+U^{*}V_{t}^{\top}-U_{t}V_{t}^{\top}-U^{*}V^{*\top}\|_{F({\Omega_{t}})}+\|S^{*}\|_{F({\Omega_{t}})}$
$=\|U_{t}(V^{*\top}-V_{t}^{\top})-U^{*}(V^{*\top}-V_{t}^{\top})\|_{F({\Omega_{t}})}+\|S^{*}\|_{F({\Omega_{t}})}$
$=\|\Delta U^{*}_{t}\Delta V_{t}^{*\top}\|_{F({\Omega_{t}})}+\|S^{*}\|_{F({\Omega_{t}})}.$
Proof of Lemma D.5.
The analogue of the objective function in eq. 1 on the set $I_{t}^{\text{out}}$ can be written as
$\mathcal{L}^{t}_{I_{t}^{\text{out}}}(U_{t+1},V_{t+1})$ $=\|U_{t}V_{t+1}^{\top}+U_{t+1}V_{t}^{\top}-U_{t}V_{t}^{\top}-L^{*}-S^{*}\|_{F(I_{t}^{\text{out}})}$
$=\|\left(U_{t}V_{t+1}^{\top}+U_{t+1}V_{t}^{\top}-U_{t}V_{t}^{\top}{-U_{t+1}V_{t+1}^{\top}}\right)+\left(U_{t+1}V_{t+1}^{\top}-L^{*}-S^{*}\right)\|_{F(I_{t}^{\text{out}})}$
$=\|\left(L_{t+1}-L^{*}-S^{*}\right)-(U_{t+1}-U_{t})(V_{t+1}-V_{t})^{\top}\|_{F(I_{t}^{\text{out}})}.$
Note that the second term is $\Delta U_{t}\Delta V_{t}^{\top}$. Hence eq. 59 follows by the triangle inequality. ∎
E.3 Proof of Lemma D.8
Proof.
We denote by $u_{i}$ and $v_{j}$ the $i$-th row and the $j$-th column of $\mathcal{P}_{\Omega}(L-L^{*})$, respectively. We denote by $u_{i}^{(k)}$ the element of $u_{i}$ with the $k$-th largest magnitude, and a similar definition for $v_{j}^{(k)}$. As $\Lambda=\text{support}\left(\mathcal{T}_{\gamma\alpha}\left(L-X,\Omega\right)\right)$, for any $(i,j)\in\Omega\setminus\Lambda$
$|(L-X)_{i,j}|\leq\max\left\{\left|(L-X)_{(i,\cdot)}^{(\lceil\gamma\alpha r_{i}\rceil)}\right|,\left|(L-X)_{(\cdot,j)}^{(\lceil\gamma\alpha c_{j})\rceil}\right|\right\}.$ (76)
By assumption 3 $S^{*}\in\mathcal{S}_{\alpha}^{\Omega}$ there are at most $\alpha r_{i}$ corrupted entries in the $i$th row. Since $\gamma>1$, the set of $\lceil\gamma\alpha r_{i}\rceil$ largest entries in the $i$th row of $\mathcal{P}_{\Omega}(L-X)$ contains at least $\lceil(\gamma-1)\alpha r_{i}\rceil$ non corrupted entries. Therefore, the $\lceil\gamma\alpha r_{i}\rceil$ largest entry in the $i$th row of $\mathcal{P}_{\Omega}(L-X)$ is smaller than the $\lceil(\gamma-1)\alpha r_{i}\rceil$ largest entry in the $i$th row of $\mathcal{P}_{\Omega}(L-L^{*})$. Formally
$|(L-X)_{(i,\cdot)}^{(\lceil\gamma\alpha r_{i}\rceil)}|\leq\left|u_{i}^{\lceil(\gamma-1)\alpha|\Omega_{(i.\cdot)}|\rceil}\right|.$ (77)
The same argument can be applied to each column $j$. Inserting eq. 77 and the analogous inequality for the $j$th column into eq. 76 yields,
$|(L-X)_{i,j}|\leq\max\left\{\left|u_{i}^{\lceil(\gamma-1)\alpha\cdot r_{i}\rceil}\right|,\left|v_{j}^{\lceil(\gamma-1)\alpha\cdot c_{j}\rceil}\right|\right\}.$ (78)
Next we upper bound $\left|u_{i}^{\lceil(\gamma-1)\alpha\cdot r_{i}\rceil}\right|$. To this end note that,
$\|u_{i}\|^{2}=\sum_{k=1}^{r_{i}}(u_{i}^{(k)})^{2}\geq\sum_{k=1}^{\lceil(\gamma-1)\alpha\cdot r_{i}\rceil}(u_{i}^{(k)})^{2}\geq\sum_{k=1}^{\lceil(\gamma-1)\alpha\cdot r_{i}\rceil}\left|u_{i}^{\lceil(\gamma-1)\alpha\cdot r_{i}\rceil}\right|^{2}\geq(\gamma-1)\alpha\cdot r_{i}\left|u_{i}^{\lceil(\gamma-1)\alpha\cdot r_{i}\rceil}\right|^{2}.$
Therefore,
$\left|u_{i}^{\lceil(\gamma-1)\alpha\cdot r_{i}\rceil}\right|^{2}\leq\frac{\|u_{i}\|^{2}}{(\gamma-1)\alpha r_{i}}.$ (79)
The same argument can be applied to $\left|v_{j}^{\lceil(\gamma-1)\alpha\cdot c_{j}\rceil}\right|$. Inserting eq. 79 into eq. 78 gives
$|(L-X)_{i,j}|^{2}\leq\frac{\|u_{i}\|^{2}}{(\gamma-1)\alpha r_{i}}+\frac{\|v_{j}\|^{2}}{(\gamma-1)\alpha c_{j}}.$
Summing over the entries in $(\Omega\setminus\Lambda)\cap\Lambda_{*}$ yields,
$\|L-X\rVert^{2}_{F({(\Omega\setminus\Lambda)\cap\Lambda_{*}})}$ $\leq\frac{1}{(\gamma-1)\alpha}\sum_{(i,j)\in{(\Omega\setminus\Lambda)\cap\Lambda_{*}}}\frac{\|u_{i}\|^{2}}{r_{i}}+\frac{\|v_{j}\|^{2}}{c_{j}}.$
Since $(\Omega\setminus\Lambda)\cap\Lambda_{*}\subseteq\Lambda_{*}$,
$\|L-X\rVert^{2}_{F({(\Omega\setminus\Lambda)\cap\Lambda_{*}})}$ $\leq\frac{1}{(\gamma-1)\alpha}\sum_{(i,j)\in{\Lambda_{*}}}\frac{\|u_{i}\|^{2}}{r_{i}}+\frac{\|v_{j}\|^{2}}{c_{j}}$
$\leq\frac{1}{(\gamma-1)\alpha}\left[\sum_{i\in[n_{1}]}\sum_{j\in{\Lambda_{*(i,\cdot)}}}\frac{\|u_{i}\|^{2}}{r_{i}}+\sum_{j\in[n_{2}]}\sum_{i\in{\Lambda_{*(\cdot,j)}}}\frac{\|v_{j}\|^{2}}{c_{j}}\right].$
By the assumption $S^{*}\in\mathcal{S}_{\alpha}^{\Omega}$, it follows that $|\Lambda_{*(i,\cdot)}|\leq\alpha r_{i}$ and $|\Lambda_{*(\cdot,j)}|\leq\alpha c_{j}$. Therefore,
$\|L-X\rVert^{2}_{F({(\Omega\setminus\Lambda)\cap\Lambda_{*}})}$ $\leq\frac{1}{(\gamma-1)\alpha}\left[\alpha\sum_{i\in[n_{1}]}{\|u_{i}\|^{2}}+\alpha\sum_{j\in[n_{2}]}{\|v_{j}\|^{2}}\right]$
$=\frac{2}{\gamma-1}\lVert\mathcal{P}_{\Omega}(L-L^{*})\rVert^{2}_{F}=\frac{2}{\gamma-1}\lVert L-L^{*}\rVert^{2}_{F(\Omega)}.$
Taking the square root on both sides completes the proof. ∎
Appendix F Proof of Lemma E.1
In our proof we would use the following lemma from Zilber and Nadler, (2022) [Lemma SM2.4].
Lemma F.1.
Let $L^{*}\in\mathbb{R}^{n_{1}\times n_{2}}$ be a matrix of rank $r$, and denote $Z^{*}=\textnormal{b-SVD}_{r}(L^{*})$. Then for any $Z=\begin{pmatrix}U\\ V\end{pmatrix}\in\mathbb{R}^{(n_{1}+n_{2})\times r}$,
$d_{P}^{2}(Z,Z^{*})\leq\frac{1}{(\sqrt{2}-1)\sigma_{r}^{*}}\left(\|UV^{\top}-L^{*}\|_{F}^{2}+\frac{1}{4}\|U^{\top}U-V^{\top}V\|_{F}^{2}\right).$ (80)
Proof of Lemma E.1.
We prove for $U_{t}$ and $U_{0}$. A similar proof follows for $V_{t}$ and $V_{0}$. By the triangle inequality
$\sigma_{1}(U_{t+1})\leq\sigma_{1}(U_{t})+\|U_{t+1}-U_{t}\|_{op}.$
Hence by induction
$\sigma_{1}(U_{t+1})\leq\sigma_{1}(U_{0})+\sum_{k=0}^{t}\|U_{k+1}-U_{k}\|_{op}\leq\sigma_{1}(U_{0})+\sum_{k=0}^{t}\|U_{k+1}-U_{k}\|_{F}.$ (81)
Since $(U_{k+1},V_{k+1})\in\mathcal{C}(U_{k},V_{k},\frac{\sigma_{r}^{*}}{4^{k+1}c_{e}\sqrt{\kappa}})$, see eq. 8,
$\|U_{k+1}-U_{k}\|_{F}\leq\sqrt{\frac{\sigma_{r}^{*}}{4^{k+1}c_{e}\sqrt{\kappa}}}=\sqrt{\frac{\sigma_{r}^{*}}{c_{e}\sqrt{\kappa}}}\frac{1}{2^{k+1}}.$
Therefore for $c_{e}\geq 1$,
$\sum_{k=0}^{t}\|U_{k+1}-U_{k}\|_{F}\leq\sqrt{\frac{\sigma_{r}^{*}}{c_{e}\sqrt{\kappa}}}\sum_{k=0}^{t}\frac{1}{2^{k+1}}\leq\sqrt{\frac{\sigma_{r}^{*}}{c_{e}\sqrt{\kappa}}}\leq\frac{\sqrt{{\sigma_{r}^{*}}}}{\sqrt[4]{\kappa}}\leq\sqrt{\sigma_{1}^{*}}.$ (82)
Next we bound $\sigma_{1}(U_{0})$. Denote $Z^{*}=\begin{pmatrix}U^{*}\\ V^{*}\end{pmatrix}=\text{b-SVD}_{r}(L^{*})$. By our assumption $(U_{0},V_{0})\in\mathcal{B}_{\textnormal{err}}\left(\frac{1}{c_{e}\sqrt{\kappa}}\right)\cap\mathcal{B}_{\textnormal{bln}}\left(\frac{1}{2c_{l}}\right)$. Therefore by Lemma F.1
$d_{P}^{2}(Z_{0},Z^{*})\leq\frac{1}{(\sqrt{2}-1)}\left(\frac{1}{c_{e}^{2}\kappa}+\frac{1}{16c_{l}^{2}}\right)\sigma_{r}^{*}\leq\sigma_{1}^{*}.$ (83)
Let $P$ be the minimizer of the Procrustes distance between $U_{0}$ and $U^{*}$. Note that $|\sigma_{1}(U_{0})-\sigma_{1}(U^{*})|=|\sigma_{1}(U_{0})-\sigma_{1}(U^{*}P)|$. By Weyl’s inequality
$|\sigma_{1}(U_{0})-\sigma_{1}(U^{*}P)|\leq\|U_{0}-U^{*}P\|_{op}\leq\|U_{0}-U^{*}P\|_{F}=d_{P}(U_{0},U^{*})\leq d_{P}(Z_{0},Z^{*}).$
Hence $|\sigma_{1}(U_{0})-\sigma_{1}(U^{*})|\leq d_{P}(Z_{0},Z^{*})$. By the triangle inequality and eq. 83
$\sigma_{1}(U_{0})\leq\sigma_{1}(U^{*})+d_{P}(Z_{0},Z^{*})\leq 2\sqrt{\sigma_{1}^{*}}.$ (84)
Inserting eq. 82 and eq. 84 into eq. 81 completes the proof. ∎